### v2

# Cross Entropy versus Label Smoothing: A Neural Collapse Perspective

Abstract
Label smoothing loss is a widely adopted technique to mitigate overfitting in deep neural networks. This paper studies label smoothing from the perspective of Neural Collapse (NC), a powerful empirical and theoretical framework which characterizes model behavior during the terminal phase of training. We first show empirically that models trained with label smoothing converge faster to neural collapse solutions and attain a stronger level of neural collapse. Additionally, we show that at the same level of NC1, models under label smoothing loss exhibit intensified NC2. These findings provide valuable insights into the performance benefits and enhanced model calibration under label smoothing loss. We then leverage the unconstrained feature model to derive closed-form solutions for the global minimizers for both loss functions and further demonstrate that models under label smoothing have a lower conditioning number and, therefore, theoretically converge faster. Our study, combining empirical evidence and theoretical results, not only provides nuanced insights into the differences between label smoothing and cross-entropy losses, but also serves as an example of how the powerful neural collapse framework can be used to improve our understanding of DNNs.
Machine Learning, ICML
1 Introduction
The effectiveness of a deep neural network (DNN) hinges significantly on the choice of the loss function during training. While cross-entropy loss is one of the most popular choices for classification tasks, many alternatives with improved empirical performance have been proposed. Among these, label smoothing (Szegedy et al., 2016; Müller et al., 2019; Lukasik et al., 2020) has emerged as a common technique to enhance the performance of DNNs. Instead of supervising the model training with one-hot key labels, label smoothing introduces a soft target label by blending the hard target label with a uniform distribution over the labels. This procedure is generally understood as a means of regularisation for improving the model’s generalizability.
In this paper, we examine the benefits of label smoothing from the perspective of neural collapse (Papyan et al., 2020), a new and powerful framework for obtaining an improved understanding of DNNs.
Background and Related Work
Many DNNs consist of a non-linear feature extractor followed by a linear classification layer. Neural collapse (Papyan et al., 2020) describes the geometric properties of the features produced by the feature extractor and the weight vectors of the classifier during the terminal phase of training:
  • (NC1): The learned features of the samples within the same class approach their respective class means.
  • (NC2): The collapsed features from different classes and the classification weight vectors approach the vertices of a simplex equiangular tight frame (ETF).
  • (NC3): Up to rescaling, the linear classifier weights approach the corresponding class means, giving a self-dual configuration.
It has been empirically verified that neural collapse occurs across a broad spectrum of neural network structures, across a variety of settings, and under different loss functions (Papyan et al., 2020; Han et al., 2021; Zhou et al., 2022b).
In addition to being observed empirically, neural collapse has also been proven to arise mathematically. Using approximation models for DNNs, such as unconstrained feature models (Mixon et al., 2020) or layer-peeled models (Fang et al., 2021), it has been demonstrated that the solutions minimizing the loss functions for these models exhibit idealized neural collapse properties NC1-NC3. These optimization models simplify a DNN by treating the last-layer features as free variables to optimize over, which is justified due to the expressiveness of DNNs (Hornik, 1991; Cybenko, 1989; Lu et al., 2017; Shaham et al., 2018). A series of studies have employed Unconstrained Feature Models (UFMs) under distinct loss functions and regularizations (Wojtowytsch et al., 2020; Zhu et al., 2021; Dang et al., 2023; Lu & Steinerberger, 2022; Tirer & Bruna, 2022; Tirer et al., 2023; Yaras et al., 2022; Zhou et al., 2022b). Importantly, these studies consistently reveal that the global minimizers of the empirical risk function under these models align with the characterization of neural collapse proposed in (Papyan et al., 2020).
Zhou et al.(Zhou et al., 2022b) established that under the UFM model, the global minimizers for a wide range of loss functions, including cross-entropy loss and label smoothing loss, have the idealized neural collapse properties and that the UFM model has a benevolent landscape allowing the global minimizer to be effectively attained using iterative algorithms. However, their investigation does not provide insights into why label smoothing consistently outperforms cross-entropy loss, or why label smoothing converges faster during training. In this paper, leveraging the neural collapse framework, we conduct an in-depth investigation of training under these loss functions, aiming to explain the reasons behind the observed superiority of label smoothing loss over cross-entropy loss.
Our Contributions
We first conduct a comprehensive empirical comparison of cross-entropy loss and label smoothing loss during the training process. Specifically, we carefully study how the last layer features and linear classifiers evolve during training. We show:
  • 1.
    Compared with CE loss, models trained under LS loss exhibit accelerated convergence in terms of training error, testing error, and neural collapse metrics. Furthermore, they converge to a more pronounced level of NC1 and NC2.
  • 2.
    Along with accelerated convergence, label smoothing loss maintains a distinct balance between NC1 and NC2. Notably, as compared with CE loss, LS loss results in a more pronounced level of NC2 when reaching a comparable level of NC1. We argue that this phenomenon originates from the implicit inductive bias introduced by label smoothing (LS), which equalizes the logits of all non-target classes, and thus promotes the emergence of a simplex ETF structure in both the learned features and the classifier vectors. We posit that the emphasis on NC2 in LS, promoting maximal separable features between classes, enhances the model’s generalization performance. Conversely, we argue an excessive focus on NC1 causes the features to overly specialize in the training data, hindering the model’s ability to generalize effectively.
  • 3.
    Through comprehensive empirical analysis, we demonstrate an excessive level of NC1 can make the model overconfident in its predictions, even when they are incorrect. Models trained under LS loss exhibit improved calibration by implicitly regularizing the weights and features during training.
Of equal importance to the empirical results, we perform a mathematical analysis of the convergence properties of the UFM models under CE and LS losses. While Zhou et al. (Zhou et al., 2022a) demonstrate that, for a broad class of loss functions, including CE and LS losses, the global minimizers exhibit neural collapse properties, the authors neither derive the exact form of these global minimizers nor thoroughly examine the landscape around them.
  1. 1.
    We first derive closed-form solutions for the global minimizers under both CE and LS loss functions, which explicitly depend on the smoothing parameter $\delta$.
  2. 2.
    Utilizing these closed-form solutions, we conduct a second-order theoretical analysis of the optimization landscape around their respective global optimizers. Within the UFM context, our mathematical analysis reveals that LS exhibits a more well-conditioned landscape around the global minimum, which facilitates the faster convergence observed in our empirical study.
This paper provides a significantly deeper understanding of why LS provides better convergence and performance than CE loss. Additionally, the paper illustrates how the powerful framework of neural collapse and its associated mathematical models can be employed to gain a more nuanced understanding of the “why” of DNNs. We hope that these results will inspire future research into the intricate interplay between neural collapse, convergence speed, and model generalizability.
2 Preliminaries
2.1 The Problem Setup
A deep neural network is comprised of two key components: a feature extractor and a linear classifier. The feature extractor $\phi_{\bm{\theta}}(\cdot)$ is a nonlinear mapping that maps the input $\bm{x}$ to the corresponding feature embedding $\bm{h}:=\phi_{\bm{\theta}}(\bm{x})\in\mathbb{R}^{d}$. Meanwhile, the linear classifier involves a weight matrix $\bm{W}=[\bm{w}_{1},\bm{w}_{2},\cdots,\bm{w}_{K}]\in\mathbb{R}^{d\times K}$ and a bias vector $\bm{b}\in\mathbb{R}^{K}$. Consequently, the architecture of a deep neural network is captured by the following equation:
$f_{\Theta}(\bm{x}):=\bm{W}^{\top}\phi_{\bm{\theta}}(\bm{x})+\bm{b},$ (1)
where $\Theta:=\{\bm{\theta},\bm{W},\bm{b}\}$ represents the set of all model parameters. In this work, we consider training a deep neural network using a balanced dataset denoted as $\{(\bm{x}_{ki},\bm{y}_{ki})\}_{1\leq k\leq K,1\leq i\leq n}$. This dataset consists of samples distributed across $K$ distinct classes, with $n$ samples allocated per class. Here, $\bm{x}_{ki}$ represents the $i$-th sample from the $k$-th class, and $\bm{y}_{ki}$ is a one-hot vector with unity solely in the $k$-th entry. Our objective is to learn the parameters $\Theta$ by minimizing the empirical risk over the total $N=nK$ training samples:
$\min_{\Theta}\frac{1}{N}\sum_{k=1}^{K}\sum_{i=1}^{n}{l}\left(f_{\Theta}(\bm{x}_{ki}),\bm{y}_{ki}\right)+\frac{\lambda}{2}\|\Theta\|^{2}_{F},$ (2)
where $l(\cdot,\cdot)$ denotes the chosen loss function, and $\lambda>0$ is the regularization parameter (i.e., the weight decay parameter).
2.2 Training Losses
To simplify the notation, we use $\bm{z}=\bm{W}^{\top}\phi_{\bm{\theta}}(\bm{x})+\bm{b}$ to represent the network’s output logit vector for a given input $\bm{x}$, and $\bm{p}=\text{Softmax}(\bm{z})$ to denote the predicted distribution from the model. The cross-entropy between the target distribution $\bm{y}$ and the predicted distribution $\bm{p}$ is defined as $l_{CE}(\bm{y},\bm{p})=-\sum_{k}y_{k}\log(p_{k})$, where $\bm{y}$ is a one-hot vector with a value of $1$ in the dimension corresponding to the target class. In contrast, LS loss minimizes the cross-entropy between the smoothed soft label $\bm{y}^{\delta}$ and the predicted distribution $\bm{p}$, denoted as $l_{CE}(\bm{y}^{\delta},\bm{p})$, where the soft label $\bm{y}^{\delta}=(1-\delta)\bm{y}+(\delta/K)\mathbf{1}_{K}$ combines the hard ground truth label $\bm{y}$ with a uniform distribution over the labels. Here $\mathbf{1}_{K}$ denotes the K-dimensional vector of all ones. The hyper-parameter $\delta$ determines the degree of smoothing.
For simplicity, we use the following formulation to represent both CE loss and LS loss:
$l_{CE}(\bm{y}^{\delta},\bm{p})=-\sum_{k}y_{k}^{\delta}\log(p_{k}),$ (3)
where $y^{\delta}_{k}=(1-\delta)y_{k}+\delta/K$. The provided loss corresponds to CE loss when $\delta=0$, and for any other value of $\delta\in(0,1)$, it represents LS loss.
3 Empirical Analysis of Cross-Entropy and Label Smoothing Losses
This section conducts a comprehensive empirical comparison between cross-entropy (CE) loss and label smoothing (LS) loss from the perspective of neural collapse. In Section 3.1, we examine the convergence of the model to neural collapse solutions under both CE and LS losses. Our results demonstrate that while models under both loss functions converge to neural collapse, LS loss induces faster convergence and reaches a more pronounced level of neural collapse. Beyond convergence rates, Section 3.2 delves into the dynamics of convergence, revealing that LS loss introduces a bias toward solutions with a symmetric simplex ETF structure, thereby enforcing NC2. In Section 3.3, we attempt to understand the impact of label smoothing on model calibration from the perspective of neural collapse. Finally, Section 3.4 further investigates how the smoothing hyperparameter $\delta$ influences neural collapses during the model training process.
Experiment Setup. We conducted experiments on CIFAR-10 (Krizhevsky et al., 2009), CIFAR-100, and STL-10 (Coates et al., 2011). For consistency with prior studies (Papyan et al., 2020; Zhu et al., 2021), we adopt ResNet-18 (He et al., 2016) as the backbone network for CIFAR-10 and ResNet-50 (He et al., 2016) as the backbone for CIFAR-100 and STL-10.
To focus on the behaviors associated with neural collapse and minimize the influence of other factors, we adopt standard preprocessing techniques without introducing any data augmentation for the training data. To comprehensively analyze the model’s behavior during the terminal phase training, we extend the training period to 800 epochs. For all datasets, we employ a default batch size of 64, stochastic gradient descent with a momentum of 0.9, weight decay of $5\times 10^{-4}$, and a learning rate that is initialized at 0.05 and undergoes multi-step decay, decreasing by a factor of 0.1 at epochs 150 and 350.
Figure 1: Neural Collapse with Cross-Entropy and Label Smoothing Losses. Columns from left to right represent the model’s misclassification error rate (Train/Test), NC1, NC2, NC3, and the average norm of the classification weight vectors/class mean features under Cross-Entropy (CE) and Label Smoothing (LS) losses.
Figure 2: The leftmost column shows a scatter diagram for NC1 vs. NC2 under CE/LS losses with colors indicating the testing error rate. The right three columns investigate the impact of the smoothing hyperparameter $\delta$. The 2nd, 3rd, and 4th columns represent (a) NC metrics vs. $\delta$, (b) Average norm of classifier vectors and mean features vs. $\delta$, and (c) Testing error rate vs. $\delta$.
Metrics for Measuring NC. We assess neural collapse (NC) in the last-layer features and the classifiers using metrics based on the properties introduced in Section 1, with metrics similar to those presented in (Papyan et al., 2020). As NC1-NC3 leads to NC4, our primary focus is on NC1-NC3. For convenience, we denote the global mean and class mean of the last-layer features as:
$\bm{h}_{G}=\frac{1}{N}\sum_{k=1}^{K}\sum_{i=1}^{n}\bm{h}_{ki},\quad\bar{\bm{h}}_{k}=\frac{1}{n}\sum_{i=1}^{n}\bm{h}_{ki},(1\leq k\leq K).$
(NC1) Within class variability measures the relative magnitude of the within-class covariance matrix $\Sigma_{W}:=\frac{1}{N}\sum_{k=1}^{K}\sum_{i=1}^{n}(\bm{h}_{ki}-\bar{\bm{h}}_{k})(\bm{h}_{ki}-\bar{\bm{h}}_{k})^{\top}$ compared to the between-class covariance matrix $\Sigma_{B}:=\frac{1}{K}\sum_{k=1}^{K}(\bar{\bm{h}}_{k}-\bm{h}_{G})(\bar{\bm{h}}_{k}-\bm{h}_{G})^{\top}$ of the last-layer features. It is formulated as:
$NC_{1}=\frac{1}{K}trace\left(\Sigma_{W}\Sigma_{B}^{\dagger}\right),$ (4)
where $\Sigma_{B}^{\dagger}$ denotes the pseudo inverse of $\Sigma_{B}$.
(NC2) Distance to simplex ETF quantifies the difference between the product of the classifier weight matrix and the centered class mean feature matrix, and a simplex ETF, defined as follows:
$NC_{2}:=\left\|\frac{{\bm{W}}^{\top}\mkern 1.5mu\overline{\mkern-1.5mu\bm{H}\mkern-1.5mu}\mkern 1.5mu}{\left\|{\bm{W}}^{\top}\mkern 1.5mu\overline{\mkern-1.5mu\bm{H}\mkern-1.5mu}\mkern 1.5mu\right\|_{F}}-\frac{1}{\sqrt{K-1}}\left(\bm{I}_{K}-\frac{1}{K}\mathbf{1}_{K}\mathbf{1}_{K}^{\top}\right)\right\|_{F},$ (5)
where $\mkern 1.5mu\overline{\mkern-1.5mu\bm{H}\mkern-1.5mu}\mkern 1.5mu=[\bar{\bm{h}}_{1}-\bm{h}_{G},\cdots,\bar{\bm{h}}_{K}-\bm{h}_{G}]\in\mathbb{R}^{d\times K}$ represents centered class mean matrix.
When the bias $\bm{b}$ is an all-zero vector or a constant vector, $NC_{2}=0$ implies that the average logit matrix, defined as $\mkern 1.5mu\overline{\mkern-1.5mu\bm{Z}\mkern-1.5mu}\mkern 1.5mu={\bm{W}}^{\top}\mkern 1.5mu\overline{\mkern-1.5mu\bm{H}\mkern-1.5mu}\mkern 1.5mu+\bm{b}\mathbf{1}_{K}^{T}$, satisfies $\mkern 1.5mu\overline{\mkern-1.5mu\bm{Z}\mkern-1.5mu}\mkern 1.5mu=a\left(\bm{I}-\frac{1}{K}\mathbf{1}_{K}\mathbf{1}_{K}^{\top}\right)$ for some constant $a$, i.e., $\mkern 1.5mu\overline{\mkern-1.5mu\bm{Z}\mkern-1.5mu}\mkern 1.5mu$ is a simplex ETF.
(NC3) Self-duality measures the distance between the classifier weight matrix $\bm{W}$ and the centered class-means $\mkern 1.5mu\overline{\mkern-1.5mu\bm{H}\mkern-1.5mu}\mkern 1.5mu$:
[…]
It is evident that when $NC_{2}$ in (5) and $NC_{3}$ in (6) both reach zero, the matrices $\bm{W}$ and $\bm{H}$ form the same simplex ETF up to some scaling. Thus our definitions of NC1-NC3 capture the same concepts as the definitions in (Papyan et al., 2020) and (Zhu et al., 2021). We say that neural collapse occurs if NC1, NC2, and NC3 collectively approach zero during the Terminal Phase of Training (TFT).
3.1 Terminal Phase Training under Label Smoothing and Cross-Entropy Loss
In this section, we compare Label Smoothing (LS) and Cross-Entropy (CE) loss during TFT. We conduct experiments on CIFAR-10, CIFAR-100, and STL10 datasets, with LS loss employing a default smoothing hyperparameter of $\delta=0.05$. Here, we only present the results for CIFAR-10 and CIFAR-100; the results for STL10 are similar and can be found in Appendix A.2.
The leftmost column in Figure 1 (with a detailed view provided in Figure 4 in the Appendix) illustrates the progression of training and testing errors throughout the training process. As expected, models trained under LS loss exhibit lower testing errors compared to those under CE loss for both datasets, showcasing an improvement in model generalizability with LS loss. Furthermore, models trained with LS loss exhibit faster convergence for both training and testing errors. Figure 1 additionally presents the three neural collapse metrics for models trained under both cross-entropy and label smoothing losses. While the values of NC3, representing the alignment of $\bm{W}$ and $\bm{H}$, remain consistently low and comparable under both loss functions, models with LS loss exhibit faster convergence in both NC1 and NC2 and eventually reach lower levels for both NC1 and NC2. We provide theoretical evidence for this empirical observation in Section 4, and in Section 3.3, we further establish a connection between NC1 and model calibration under both loss functions.
3.2 Label Smoothing Induces Enhanced NC2
Along with the faster convergence to neural collapse under LS loss, we further observe that LS loss maintains a distinct balance between NC1 and NC2 throughout training. Specifically, we find that for the same level of NC1, LS consistently provides a more pronounced manifestation of NC2 compared to CE Loss. The leftmost column of Figure 2 provides a scatter plot of NC1 and NC2 under both loss functions, with the NC metrics recorded every $10$ training epochs. Across all three datasets (with the plot for STL10 provided in Figure 6), the data points under LS loss consistently position to the left of those for CE loss. This observation suggests that at equivalent levels of NC1, LS loss induces an intensified level of NC2.
To gain insight into this phenomenon, we now closely examine the formulations of cross-entropy and label smoothing loss. Given an input $\bm{x}$, the output logit and the predicted distribution are equal to $\bm{z}=f_{\Theta}(\bm{x})$ and $\bm{p}=\text{Softmax}(\bm{z})$, respectively. As per Equation 3, assuming the observation $\bm{x}$ is from the $k$-th class, CE loss is formulated as $l_{CE}=-\log(p_{k})$. To minimize cross-entropy loss, the emphasis is solely on making the logit of the target class larger than the logits of non-target classes without constraining the logit variation among the non-target classes.
Figure 3: Left-most plot: Comparison of test error and entropy for correctly and incorrectly classified test samples. Middle plot: NC1 for the training set, testing set, correctly classified testing samples, and incorrectly classified samples. Right plot: Model classification error rate and the test Expected Calibration Error (ECE).
On the other hand, LS with smoothing hyperparameter $\delta$ is given by
$l_{LS}=-\left(\left(1-\frac{K-1}{K}\delta\right)\log(p_{k})+\frac{\delta}{K}\sum_{l\neq k}\log(p_{l})\right).$
According to Jensen’s inequality, we have
$\frac{1}{K-1}\sum_{l\neq k}\log(p_{l})\leq\log\left(\frac{1}{K-1}\sum_{l\neq k}p_{l}\right)=\log\left(1-p_{k}\right),$
with equality achieved only if $p_{l}=p_{l^{\prime}}$ (for $l,l^{\prime}\neq k$). This implies that LS loss reaches its minimum only if the predicted probabilities for non-target classes are all equal. Thus, LS loss implicitly includes an inductive bias towards equalizing the logits of the non-target classes, which is not present in CE loss. This property significantly aligns with the definition of $NC_{2}$ in (5). As demonstrated in Section A.1 in the appendix, for deep neural networks with L2 regularization, the convergence of $NC_{2}$ as defined in (5) to zero indicates that both the classification vectors $\bm{W}$ and class mean features $\mkern 1.5mu\overline{\mkern-1.5mu\bm{H}\mkern-1.5mu}\mkern 1.5mu$ converge to a simplex ETF structure. Thus, label smoothing loss inherently carries an inductive bias towards reinforcing NC2.
The convergence of classifier weights and class mean features towards simplex ETF promotes the development of maximally separable features and classifiers, a critical factor for enhancing overall model performance. Conversely, an excessive level of NC1 suggests potential overfitting of the model to the training data, posing a risk to its generalizability. We posit the improvement in model generalizability observed with LS loss can be attributed to the prioritization of enhancing NC2 over NC1 during the training process.
3.3 Label Smoothing and Model Calibration
Previous studies (Müller et al., 2019) have emphasized the effectiveness of label smoothing in improving model calibration. We now explore the connection between model calibration and neural collapse.
DNN models often suffer from poor calibration, where the assigned probability values to class labels tend to overestimate the actual likelihood of the correctness. This issue arises from the high capacity of DNN models, making them prone to overfitting the training data. Guo et al. (Guo et al., 2017) proposed the expected calibration error (ECE) as a measure of model calibration. Additionally, they proposed temperature scaling as an effective post-processing method for model calibration. In this section, we delve into the nuanced implications of NC1 and the norms of both $\bm{W}$ and $\bm{H}$ on model calibration, which helps to understand the distinctions in model calibration under CE and LS loss functions.
Excessive NC1 adversely impacts model calibration. To investigate the impact of NC1 on model calibration, we conduct experiments on CIFAR-10 to examine how miscalibration occurs during training. The left plot of Figure 3 shows the average CE loss for the test set, along with the average CE loss for correctly/incorrectly classified test samples. Additionally, it presents the average entropy of the predicted probabilities for the corresponding groups. The middle of Figure 3 presents the NC1 metric for the test data, as well as the NC1 metric for correctly/incorrectly classified test samples. The right column of Figure 3 shows the training and testing misclassification errors along with the test set ECE.
The left plot in Figure 3 indicates that while the overall testing loss has a consistent decreasing trend, there is a noteworthy increase in the average loss for incorrectly classified test samples after 350 epochs. Additionally, there is a consistent reduction in the entropies of the predicted probabilities for both correctly and incorrectly classified test samples throughout training. Further, the middle plot in Figure 3 indicates that NC1 for both correctly and incorrectly classified test samples consistently decreases during training. Note that for misclassified testing samples, we measure the within-class variance with respect to the incorrectly predicted class. Therefore, a smaller NC1 suggests that their feature vectors align more closely with the mean of the incorrectly predicted class, making the model more confident about its incorrect predictions. This results in a smaller entropy and contributes to miscalibration, as evidenced by the right plot of Figure 3, where the test ECE starts to rise post 350 epochs, indicating that the model becomes miscalibrated. In conclusion, we assert that an excessive level of NC1 causes the model to become overly confident in its predictions, even when those predictions are incorrect, thereby adversely impacting model calibration.
Weight and feature regularization improve model calibration. By increasing the norm of the classification weight vectors or the feature embeddings, we amplify the logits $\bm{Z}$, thereby increasing the confidence of the network’s predictions. Notably, the cross-entropy loss is minimized when $\|\bm{Z}\|\rightarrow\infty$. In contrast, models trained with LS loss exhibit significantly smaller norms for the classification weight vectors and the last-layer features, as illustrated in the rightmost column of Figure 1. The lower norms effectively mitigate the overconfidence issue in the model.
Temperature scaling as an effective post-processing method to improve model calibration (Guo et al., 2017), involves dividing a network’s logits by a scalar $T>0$ before applying softmax, thereby softening (for $T>1$) or sharpening (for $T<1$) the predicted probabilities. It is noteworthy that adjusting $\delta$ in LS loss can achieve improved model calibration without the need for temperature scaling. The smoothing parameter $\delta$ controls the regularization effects on $\bm{W}$ and $\bm{H}$, as empirically illustrated in Figure 2 and theoretically demonstrated in Theorem 4.1. For further evidence of this claim, in Section A.3 of the Appendix, we present a 20-bin reliability plot (Niculescu-Mizil & Caruana, 2005) for models trained under CE and LS loss, both before and after temperature scaling for CIFAR-10.
Integrating these findings from the preceding discussions, we now address model calibration under CE and LS losses. The quantitative analysis in Table 1 presents 20-bin ECE (%) values before and after temperature scaling for both loss functions, with LS loss utilizing a default smoothing parameter of 0.05. Prior to temperature scaling, the model with LS loss exhibits a lower ECE for both CIFAR-10 and CIFAR-100, due to its implicit regularization effect on $\bm{W}$ and $\bm{H}$. For STL10, the model with LS loss exhibits higher ECE due to the exceptionally low norm of its $\bm{W}$ and $\bm{H}$. Applying temperature scaling with $T=0.6$ can help magnify the norm and optimize the test ECE. However, post-temperature scaling, the model with LS loss exhibits higher ECE for all three datasets. This is because LS loss results in stronger NC1. The excessive presence of NC1, especially in testing data, leads to increased confidence in incorrectly classified test samples, detrimentally impacting model calibration.
Dataset Cross-Entropy LS-0.05
Pre T Post T Pre T Post T
CIFAR-10 5.13 1.41 (1.4) 3.78 1.89 (0.9)
CIFAR-100 9.89 5.56 (1.3) 6.73 6.52 (1.1)
STL-10 10.51 1.30 (1.5) 12.51 2.12 (0.6)
Table 1: ECE (%) computed for models trained with CE and LS loss, both before and after temperature scaling. The optimal temperature for each model is indicated in parentheses.
3.4 Impact of the Smoothing Hyperparameter
In Section 3.1 and 3.2, a default smoothing hyperparameter of $\delta=0.05$ is utilized. This section explores how the choice of $\delta$ impacts the model’s convergence to neural collapse and its generalizability. Specifically, we consider various values for $\delta$ within the interval $[0,1)$.
Figure 2 presents the results of the experiments. The experiments yield several important insights. Firstly, the smoothing hyperparameter significantly influences the model’s convergence to neural collapse. The curves of NC1 and NC2 as a function of $\delta$ (2nd column in Figure 2) reveal a U-shaped trend, with the levels of NC3 remaining relatively consistent across different values of $\delta$. Exceptionally small and large values of $\delta$ result in reduced collapse for both NC1 and NC2.
Secondly, the average norms of the classifier vectors and the class mean features decrease as $\delta$ increases. This observation aligns with intuition: given features and classifier vectors that form a simplex ETF structure, decreasing their norms softens the output probabilities. For LS loss, a higher smoothing parameter $\delta$ corresponds to smoother target labels, and consequently, the features and classification vectors with lower norms can achieve close-to-zero label smoothing loss. This observation was also supported by the closed-form expressions for $\bm{W}$ and $\bm{H}$ provided in Theorem 4.1. Thirdly, the trend of the test error also exhibits a U-shape, underscoring the necessity of selecting an appropriate $\delta$. When $\delta$ approaches $1$, the nearly uniform smoothed labels do not provide effective training signals to update the parameters. Additionally, the norm of both classification vectors and the last-layer features decrease towards zero, causing the classifier to be dominated by noise, leading to rapid deterioration in model performance. Finally, we observe a surprising robustness in both neural collapse metrics and model performance for the choice of $\delta\in[0.02,0.8]$. For classification tasks with a larger number of classes, such as CIFAR-100, opting for a relatively large $\delta$ is viable.
4 Theoretical Analysis
4.1 Unconstrained Feature Model
Analyzing deep neural networks poses significant challenges because of the non-linearities and complex interactions between layers in the feature mapping $\bm{h}:=\phi_{\bm{\theta}}(\bm{x})$. The Unconstrained Feature Model (UFM) (Fang et al., 2021) simplifies DNN models by treating the features of the last layer as free optimization variables. This choice is motivated by the idea that over-parameterized DNN models are able to approximate any continuous functions (Hornik, 1991; Cybenko, 1989; Lu et al., 2017; Shaham et al., 2018).
Recall that the dimension of a feature vector $\bm{h}$ is denoted by $d$. Let $\bm{H}=\{\bm{h}_{ki}\}_{1\leq k\leq K,1\leq i\leq n}\in\mathbb{R}^{d\times N}$ be the feature matrix of all training samples with $\bm{h}_{ki}$ denoting the $\left((k-1)n+i\right)$-th column of $\bm{H}$. Under the UFM, we investigate regularized empirical risk minimization, a variant of the formulation in Equation 2:
$\min_{\bm{W},\bm{H},\bm{b}}$ $\frac{1}{N}\sum_{k=1}^{K}\sum_{i=1}^{n}{l_{CE}}(\bm{W}^{\top}\bm{h}_{ki}+\bm{b},\bm{y}_{k}^{\delta})$
$+\frac{\lambda_{W}}{2}\|\bm{W}\|^{2}_{F}+\frac{\lambda_{H}}{2}\|\bm{H}\|^{2}_{F}+\frac{\lambda_{b}}{2}\|\bm{b}\|^{2},$ (7)
where $\bm{y}_{k}^{\delta}=(1-\delta)\bm{e}_{k}+{\delta}\frac{\mathbf{1}_{K}}{K}$ is the soft label for class $k$ with a smoothing parameter $\delta$, and $\lambda_{W},\lambda_{H},\lambda_{b}>0$ are the regularization parameters.
Let $\bm{Y}=[\bm{e}_{1}\mathbf{1}_{n}^{\top},\cdots,\bm{e}_{K}\mathbf{1}_{n}^{\top}]\in\mathbb{R}^{K\times N}$ represent the matrix form of the (hard) ground truth labels. Consequently, the matrix form of the soft labels can be represented as $\bm{Y}^{\delta}=(1-\delta)\bm{Y}+\frac{\delta}{K}\mathbf{1}_{K}\mathbf{1}_{N}^{\top}$. The empirical risk minimization in (4.1) has an equivalent matrix form:
$\min_{\bm{W},\bm{H},\bm{b}}\mathcal{L}(\bm{W},$ $\bm{H},\bm{b}):=\frac{1}{N}{l_{CE}}(\bm{W}^{\top}\bm{H}+\bm{b}\mathbf{1}_{N}^{\top},\bm{Y}^{\delta})+$
$\frac{\lambda_{W}}{2}\|\bm{W}\|^{2}_{F}+\frac{\lambda_{H}}{2}\|\bm{H}\|^{2}_{F}+\frac{\lambda_{b}}{2}\|\bm{b}\|^{2},$ (8)
where ${l_{CE}}(\bm{W}^{\top}\bm{H}+\bm{b}\mathbf{1}_{N}^{\top},\bm{Y}^{\delta})$ computes the cross-entropy column-wise and takes the sum.
4.2 Theoretical Results
Within the UFM framework, Zhou et al. (Zhou et al., 2022b) demonstrate that, under a wide range of loss functions, including CE loss and LS loss, all the global minimizers satisfy neural collapses properties, and in particular $\bm{H}$ and $\bm{W}$ form aligned simplex ETFs. Moreover, they show that every critical point is either a global minimizer or a strict saddle point, which implies that these global minimizers can be effectively attained through iterative algorithms. However, their work does not provide the exact expression of the global minimizers, nor does it provide a landscape analysis (condition number) across different loss functions, which is critical for understanding the model’s convergence behavior. In contrast, our work derives the exact solutions of both $\bm{W}$ and $\bm{H}$ for UFM models with CE and LS losses. Furthermore, based on these solutions, we employ conditioning number analysis to closely compare the optimization landscape of the models surrounding their respective global minimizers, and thereby provide an explanation for the accelerated convergence observed under LS loss. The proofs of our theorems can be found in the Appendix.
Let $\lambda_{Z}:=\sqrt{\lambda_{W}\lambda_{H}}$ and let $a^{\delta}$ be defined by
  • (i)
    if $\sqrt{KN}\lambda_{Z}+\delta\geq 1$, then $a^{\delta}=0$;
  • (ii)
    if $\sqrt{KN}\lambda_{Z}+\delta<1$, then
    $a^{\delta}=\frac{1}{K}\log\left(\frac{K}{\sqrt{KN}\lambda_{Z}+\delta}-K+1)\right).$
Theorem 4.1.
(Global Optimizer). Assume that the feature dimension $d$ is no smaller than the number of classes $K$, i.e., $d\geq K$, and the dataset is balanced. Then any global optimizer of $(\bm{W},\bm{H},\bm{b})$ of (4.1) satisfies: There is partial orthogonal matrix $\bm{P}\in\mathbb{R}^{d\times K}$ (i.e., $\bm{P}^{\top}\bm{P}=\bm{I}_{K}$) such that
$(i)$ The classification weight matrix $\bm{W}$ is given by
$\bm{W}=\left(\frac{n\lambda_{H}}{\lambda_{W}}\right)^{1/4}\sqrt{a^{\delta}}\bm{P}\left(K\bm{I}_{K}-\mathbf{1}_{K}\mathbf{1}_{K}^{\top}\right).$ (9)
(ii) The matrix of last-layer features $\bm{H}$ is given by
$\bm{H}=\left(\frac{\lambda_{W}}{n\lambda_{H}}\right)^{1/4}\sqrt{a^{\delta}}\bm{P}\left(K\bm{I}_{K}-\mathbf{1}_{K}\mathbf{1}_{K}^{\top}\right)\bm{Y}.$ (10)
$(iii)$ The bias $\bm{b}$ is a zero vector, i.e. $\bm{b}=\bm{0}$.
The above theorem provides an explicit closed-form solution for the global minimizer $(\bm{W},\bm{H},\bm{b})$. Notably, our findings highlight an inverse relationship between the norm of $\bm{W}$ and $\bm{H}$ and the smoothing parameter $\delta$. This observation aligns closely with the empirical results detailed in Section 3.4.
Recalling our observations from Section 3, where LS loss demonstrates accelerated convergence in terms of NC1, NC2, and training/testing errors, we now analyze the optimization landscape under the UFM to partially explain this phenomenon. Note that the condition number of the Hessian matrix, representing the ratio of the largest to the smallest eigenvalue, plays a crucial role in convergence rate analysis (Nocedal & Wright, 1999; Trefethen & Bau, 2022). In the vicinity of a local minimizer, a smaller condition number typically signifies a faster convergence rate. We now present our main theoretical result, which, to the best of our knowledge, is the first result providing insight into the convergence rate to the optimal solution under UFM models.
Theorem 4.2.
(Optimization Landscape). Assuming a feature dimension $d$ greater than or equal to the number of classes $K$ ($d\geq K$) and a balanced dataset, we examine the Hessian of the empirical loss with respect to ${\bm{W}}$ when ${\bm{H}}$ is fixed (also the one w.r.t. ${\bm{H}}$ when ${\bm{W}}$ is fixed) at the global minimizer under both cross-entropy and label smoothing losses. We demonstrate that the Hessians corresponding to label smoothing loss consistently exhibit a smaller condition number compared to cross-entropy loss.
This theorem suggests that the LS loss contributes to a more favorable optimization landscape, characterized by a smaller condition number of the corresponding Hessian matrix. This explains why the model trained under LS loss exhibits better convergence behavior as observed in Section 3.1.
5 Discussion
We conducted a comprehensive empirical comparison of CE loss and LS loss during the training process. We found that models trained under LS loss exhibit accelerated convergence in terms of training/testing error and neural collapse metrics. Furthermore, they converge to a more pronounced level of NC1 and NC2. Along with the accelerated convergence, we found that label smoothing maintains a distinct balance between NC1 and NC2. We posit that the emphasis on NC2 in LS enhances the model’s generalization performance. Conversely, we argue an excessive focus on NC1 causes the features to overfit in the training data.
We performed a mathematical analysis of the convergence properties of the UFM models under CE and LS losses. We first derived closed-form solutions for the global minimizers under both loss functions. Then we conducted a second-order theoretical analysis of the optimization landscape around their respective global optimizers, which reveals that LS exhibits a better-conditioned optimization landscape around the global minimum, which facilitates the faster convergence observed in our empirical study.
This paper provides a significantly deeper understanding of why LS provides better convergence and performance than CE loss. Additionally, it illustrates how the powerful framework of neural collapse and its associated mathematical models can be employed to gain a more nuanced understanding of the ”why” of DNNs. We expect that these results will inspire future research into the interplay between neural collapse, convergence speed, and model generalizability.
Appendix A Additional Experiments
A.1 Metrics for Measuring NC
We assess neural collapse (NC) in the last-layer features and the classifiers using metrics based on the properties introduced in Section 1, with metrics similar to those presented in (Papyan et al., 2020). As NC1-NC3 leads to NC4, our primary focus is on NC1-NC3. For convenience, we denote the global mean and class mean of the last-layer features as:
$\bm{h}_{G}=\frac{1}{N}\sum_{k=1}^{K}\sum_{i=1}^{n}\bm{h}_{ki},\quad\bar{\bm{h}}_{k}=\frac{1}{n}\sum_{i=1}^{n}\bm{h}_{ki},(1\leq k\leq K).$
(NC1) Within class variability collapse measures the relative magnitude of the within-class covariance $\Sigma_{W}:=\frac{1}{N}\sum_{k=1}^{K}\sum_{i=1}^{n}(\bm{h}_{ki}-\bar{\bm{h}}_{k})(\bm{h}_{ki}-\bar{\bm{h}}_{k})^{\top}$ compared to the between-class covariance matrix $\Sigma_{B}:=\frac{1}{K}\sum_{k=1}^{K}(\bar{\bm{h}}_{k}-\bm{h}_{G})(\bar{\bm{h}}_{k}-\bm{h}_{G})^{\top}$ of the last-layer features. It is formulated as:
$NC_{1}=\frac{1}{K}trace\left(\Sigma_{W}\Sigma_{B}^{\dagger}\right),$
where $\Sigma_{B}^{\dagger}$ denotes the pseudo inverse of $\Sigma_{B}$.
(NC2) Convergence to simplex ETF quantifies the difference between the normalized classifier weight matrix and the centered class mean features in comparison to a normalized simplex ETF, defined as follows:
$NC_{2}:=\left\|\frac{{\bm{W}}^{\top}\mkern 1.5mu\overline{\mkern-1.5mu\bm{H}\mkern-1.5mu}\mkern 1.5mu}{\left\|{\bm{W}}^{\top}\mkern 1.5mu\overline{\mkern-1.5mu\bm{H}\mkern-1.5mu}\mkern 1.5mu\right\|_{F}}-\frac{1}{\sqrt{K-1}}\left(\bm{I}_{K}-\frac{1}{K}\mathbf{1}_{K}\mathbf{1}_{K}^{\top}\right)\right\|_{F},$ (11)
where $\mkern 1.5mu\overline{\mkern-1.5mu\bm{H}\mkern-1.5mu}\mkern 1.5mu=[\bar{\bm{h}}_{1}-\bm{h}_{G},\cdots,\bar{\bm{h}}_{K}-\bm{h}_{G}]\in\mathbb{R}^{d\times K}$ represents centered class mean matrix.
When the bias $\bm{b}$ is an all-zero vector or a constant vector, $NC_{2}$ as defined in (11) approaching zero indicates that the average logit matrix, formulated as $\mkern 1.5mu\overline{\mkern-1.5mu\bm{Z}\mkern-1.5mu}\mkern 1.5mu={\bm{W}}^{\top}\mkern 1.5mu\overline{\mkern-1.5mu\bm{H}\mkern-1.5mu}\mkern 1.5mu+\bm{b}\mathbf{1}_{K}^{T}$, satisfies the condition $\mkern 1.5mu\overline{\mkern-1.5mu\bm{Z}\mkern-1.5mu}\mkern 1.5mu=a\left(\bm{I}-\frac{1}{K}\mathbf{1}_{K}\mathbf{1}_{K}^{\top}\right)$ for some constant $a$. Remarkably, this matrix aligns with the simplex-encoding label (SEL) matrix introduced in (Thrampoulidis et al., 2022), up to a scaling factor $a$.
From Proposition B.4, we have
$\min_{\bm{W}^{\top}{\bm{H}}={\bm{Z}}}\frac{1}{2}\left(\lambda_{W}\|\bm{W}\|^{2}_{F}+\lambda_{H}\|\bm{H}\|^{2}_{F}\right)\geq\|\mkern 1.5mu\overline{\mkern-1.5mu\bm{Z}\mkern-1.5mu}\mkern 1.5mu\|_{\ast}$
where $\|\mkern 1.5mu\overline{\mkern-1.5mu\bm{Z}\mkern-1.5mu}\mkern 1.5mu\|_{\ast}$ represents the nuclear norm of $\mkern 1.5mu\overline{\mkern-1.5mu\bm{Z}\mkern-1.5mu}\mkern 1.5mu$ and the equality holds only when $\bm{H}=\mkern 1.5mu\overline{\mkern-1.5mu\bm{H}\mkern-1.5mu}\mkern 1.5mu\bm{Y}$ and $\mkern 1.5mu\overline{\mkern-1.5mu\bm{H}\mkern-1.5mu}\mkern 1.5mu=\sqrt{{\lambda_{W}}/{n\lambda_{H}}}\bm{W}$, indicating the self-duality of the class mean feature and the classification vector. Consequently, during the model training with an L2 penalty on both $\bm{W}$ and $\bm{H}$ (with the norm of $\bm{H}$ implicitly penalized by penalizing the model parameters $\Theta$), the convergence of $NC_{2}$ as defined in (5) to zero indicates the simultaneous convergence of both $\bm{W}$ and $\bm{H}$ towards the simplex ETF structure.
(NC3) Convergence to self-duality measures the distance between the classifier weight matrix $\bm{W}$ and the centered class-means $\mkern 1.5mu\overline{\mkern-1.5mu\bm{H}\mkern-1.5mu}\mkern 1.5mu$:
[…]
A.2 Neural Collapses under Cross-Entropy Loss and Label Smoothing Loss
This section presents additional visualizations supporting the conclusions drawn in the main paper. Figure 4 offers a detailed zoom-in plot illustrating the progression of training and testing errors throughout the model training process under both cross-entropy and label smoothing losses. Notably, the visual evidence suggests that models trained with label smoothing exhibit a more rapid convergence in both training and testing errors compared to their cross-entropy counterparts.
Figure 5 demonstrates how the neural collapse metrics evolve under both cross-entropy and label smoothing losses for the STL10 dataset. Importantly, the results are consistent with those observed for CIFAR-10 and CIFAR-100—models under label smoothing loss converge faster to neural collapse solutions in terms of both NC1 and NC2, achieving a stronger level of NC1 and NC2. Moreover, Figure 6 includes a scatter plot of NC2 versus NC1 throughout the model training process, accompanied by a visualization illustrating the impact of the smoothing hyperparameter $\delta$ on the training/testing loss and neural collapse (NC) metrics for the STL10 dataset.
Figure 4: Comparison of training (top row) and testing (bottom row) error rates using cross-entropy (CE) and label smoothing (LS) losses.
Figure 5: Neural Collapse Analysis for STL10. Figures from left to right represent the model’s misclassification error rate (Train/Test), NC1, NC2, NC3, and the average norm of the classification weight vectors/class mean features under cross-entropy (CE) and label smoothing (LS) losses.
Figure 6: The leftmost column visualizes NC1 vs. NC2 under CE/LS losses with colors indicating the testing error rate. The right three columns investigate the impact of the smoothing hyperparameter $\delta$. The 2nd, 3rd, and 4th columns represent (a) NC metrics vs. $\delta$, (b) Average norm of classifier vectors and mean features vs. $\delta$, and (c) Testing performance (Test error rate) vs. $\delta$.
A.3 Weight and Feature Regularization Improve Model Calibration
Figure 7: Reliability plots with 20 bins (top row) and the percentage of samples in each bin (bottom row) for models trained with cross-entropy (CE) and label smoothing (LS) losses, both pre- (left column) and post- (right column) temperature scaling. The results are based on the CIFAR-10 dataset.
Figure 7 presents a 20-bin reliability plot (Niculescu-Mizil & Caruana, 2005) for models trained under CE and LS loss for CIFAR-10. The default smoothing hyperparameter of 0.05 is applied. Additionally, we include a plot showing the percentage of samples in each confidence bin. Clearly, LS loss exhibits less confidence in its predictions and better model calibration compared to CE loss. In the right column of Figure 7, we employ temperature scaling to calibrate the model, with the hyperparameter $T$ selected through cross-validation. Notably, after temperature scaling, models under CE loss achieve a similar or even better calibration than LS loss. Leveraging label smoothing loss with a properly tuned smoothing hyperparameter serves as an implicit regularization method for the norms of the classification weight matrix $\bm{W}$ and the last-layer features $\bm{H}$ (as illustrated in Section 3.3). This approach leads to enhanced model calibration without the need for temperature scaling.
Appendix B Proofs
The first key observation of the phenomenon of neural collapse, that is (NC1), refers to a type of collapse that involves the convergence of the feature of samples from the same class to a unique mean feature vector. The second key observation of neural collapse, namely (NC2), involves these unique mean feature vectors (after recentering by their global mean) as they form an equiangular tight frame (ETF), i.e., they share the same pairwise angles and length. Before providing the theoretical poof, we first formally define the rank $K$ canonical simplex ETF in the definition below.
B.1 Basics
Definition B.1.
(K-simplex ETF) A K-simplex ETF is a collection of points in $\mathbb{R}^{d}$ specified by the columns of the matrix
$\bm{M}=\sqrt{\frac{K}{K-1}}\bm{P}\left(\bm{I}_{K}-\frac{1}{K}\mathbf{1}_{K}\mathbf{1}_{K}^{\top}\right),$
where $\bm{I}_{K}\in\mathbb{R}^{K\times K}$ is the identity matrix and $\mathbf{1}_{K}\in\mathbb{R}^{K}$ is the ones vector, and $\bm{P}\in\mathbb{R}^{d\times K}(d\geq K)$ is a partial-orthogonal matrix such that $\bm{P}^{\top}\bm{P}=\bm{I}_{K}$.
Note that the matrix $\bm{M}$ satisfies:
$\bm{M}^{\top}\bm{M}=\frac{K}{K-1}\left(\bm{I}_{K}-\frac{1}{K}\mathbf{1}_{K}\mathbf{1}_{K}^{\top}\right).$
Next, we prove a series of lemmas that will prove crucial upon establishing our main theorems.
Lemma B.2.
(Young’s Inequality) Let $p,q$ be positive numbers satisfying $\frac{1}{p}+\frac{1}{q}=1$. Then for any $a,b\in\mathbb{R}$, we have
$|ab|\leq\frac{|a|^{p}}{p}+\frac{|b|^{q}}{q},$
where the equality holds if and only if $|a|^{p}=|b|^{q}$. The case for $p=q=2$ is just the AM-GM inequality which is $|ab|\leq\frac{1}{2}\left(a^{2}+b^{2}\right)$, where the equality holds if and only if $|a|=|b|$
Lemma B.3.
For any fixed $\bm{Z}\in\mathbb{R}^{K\times N}$ and $\alpha>0$, we have
$\min_{\bm{Z}=\bm{W}^{\top}\bm{H}}\frac{1}{2\sqrt{\alpha}}\left(\|\bm{W}\|^{2}_{F}+\alpha\|\bm{H}\|^{2}_{F}\right)=\|\bm{Z}\|_{\ast}.$ (13)
Here $\|\bm{Z}\|_{\ast}$ denotes the nuclear norm of $\bm{Z}$:
$\|\bm{Z}\|_{\ast}:=\sum_{k}\sigma_{k}(\bm{Z})=trace(\Sigma),\text{ with }\bm{Z}=\bm{U}\Sigma\bm{V}^{\top},$
where $\{\sigma_{k}\}_{k=1}^{min(K,N)}$ denote the singular values of $\bm{Z}$, and $\bm{Z}=\bm{U}\Sigma\bm{V}^{\top}$ is the singular value decomposition (SVD) of $\bm{Z}$.
Proof.
Let $\bm{Z}=\bm{U}\Sigma\bm{V}^{\top}$ be the SVD of $\bm{Z}$. From the fact that $\bm{U}\bm{U}^{\top}=\mathbf{I}$, $\bm{V}\bm{V}^{\top}=\mathbf{I}$, and $trace\left(\bm{A}^{\top}\bm{A}\right)=\|\bm{A}\|^{2}_{F}$, we have
$\|\bm{Z}\|_{\ast}=trace(\Sigma)$ $=\frac{1}{2\sqrt{\alpha}}trace\left(\sqrt{\alpha}\bm{U}^{\top}\bm{U}\Sigma\right)+\frac{\sqrt{\alpha}}{2}trace\left(\frac{1}{\sqrt{\alpha}}\Sigma\bm{V}^{\top}\bm{V}\right)$
$=\frac{1}{2\sqrt{\alpha}}\left(\left\|\alpha^{1/4}\bm{U}\Sigma^{1/2}\right\|^{2}_{F}+\alpha\left\|\alpha^{-1/4}\Sigma^{1/2}\bm{V}^{\top}\right\|^{2}_{F}\right).$
This implies that there exists some $\bm{W}=\alpha^{1/4}\Sigma^{1/2}\bm{U}^{\top}$ and $\bm{H}=\alpha^{-1/4}\Sigma^{1/2}\bm{V}^{\top}$, such that $\|\bm{Z}\|_{\ast}=\frac{1}{2\sqrt{\alpha}}\left(\|\bm{W}\|^{2}_{F}+\alpha\|\bm{H}\|^{2}_{F}\right)$, which further indicates that
$\|\bm{Z}\|_{\ast}\geq\min_{\bm{Z}=\bm{W}^{\top}\bm{H}}\frac{1}{2\sqrt{\alpha}}\left(\|\bm{W}\|^{2}_{F}+\alpha\|\bm{H}\|^{2}_{F}\right).$ (14)
On the other hand, for any $\bm{W}^{\top}\bm{H}=\bm{Z}$, we have
$\|\bm{Z}\|_{\ast}$ $=trace(\Sigma)=trace(\bm{U}^{\top}\bm{Z}V)=trace(\bm{U}^{\top}\bm{W}^{\top}\bm{H}\bm{V})$
$\leq\frac{1}{2\sqrt{\alpha}}\left\|\bm{U}^{\top}\bm{W}^{\top}\right\|^{2}_{F}+\frac{\sqrt{\alpha}}{2}\left\|\bm{H}\bm{V}\right\|^{2}_{F}=\frac{1}{2\sqrt{\alpha}}\left(\|\bm{W}\|^{2}_{F}+\alpha\|\bm{H}\|^{2}_{F}\right),$
where the first inequality is guaranteed by Young’s inequality in Lemma B.2, and equality only holds when $\bm{W}\bm{U}=\sqrt{\alpha}\bm{H}\bm{V}$. The last equality follows because $\bm{U}\bm{U}^{\top}=\bm{I}$ and $\bm{V}\bm{V}^{\top}=\bm{I}$ Therefore, we have
$\|\bm{Z}\|_{\ast}\leq\min_{\bm{Z}=\bm{W}^{\top}\bm{H}}\frac{1}{2\sqrt{\alpha}}\left(\|\bm{W}\|^{2}_{F}+\alpha\|\bm{H}\|^{2}_{F}\right).$ (15)
Combining the results in (14) and (15), we complete the proof.
Proposition B.4.
Consider matrics $\bm{H}=[\bm{H}_{1},\cdots,\bm{H}_{n}]\in\mathbb{R}^{d\times N}$ and $\bm{Z}=[\bm{Z}_{1},\cdots,\bm{Z}_{n}]\in\mathbb{R}^{K\times N}$, where $\bm{H}_{i}\in\mathbb{R}^{d\times K}$ and $\bm{Z}_{i}\in\mathbb{R}^{K\times K}$ with $N=nK$. Let $\mkern 1.5mu\overline{\mkern-1.5mu\bm{H}\mkern-1.5mu}\mkern 1.5mu=\frac{1}{n}\sum_{i}\bm{H}_{i}$ and $\mkern 1.5mu\overline{\mkern-1.5mu\bm{Z}\mkern-1.5mu}\mkern 1.5mu=\frac{1}{n}\sum_{i}\bm{Z}_{i}$. Then $\bm{Z}=\bm{W}^{\top}\bm{H}$ indicates $\mkern 1.5mu\overline{\mkern-1.5mu\bm{Z}\mkern-1.5mu}\mkern 1.5mu=\bm{W}^{\top}\mkern 1.5mu\overline{\mkern-1.5mu\bm{H}\mkern-1.5mu}\mkern 1.5mu$. If $\mkern 1.5mu\overline{\mkern-1.5mu\bm{Z}\mkern-1.5mu}\mkern 1.5mu$ is a symmetric matrix, then we have
[…]
with the minimum is reached only if $\bm{H}_{i}=\mkern 1.5mu\overline{\mkern-1.5mu\bm{H}\mkern-1.5mu}\mkern 1.5mu$ (for $\forall i=1,\cdots,n$) and $\bm{W}=\sqrt{\alpha n}\mkern 1.5mu\overline{\mkern-1.5mu\bm{H}\mkern-1.5mu}\mkern 1.5mu$.
Proof.
[…]
Utilizing Lemma B.3, we further deduce:
$\min_{\mkern 1.5mu\overline{\mkern-1.5mu\bm{Z}\mkern-1.5mu}\mkern 1.5mu=\bm{W}^{\top}\mkern 1.5mu\overline{\mkern-1.5mu\bm{H}\mkern-1.5mu}\mkern 1.5mu}\frac{1}{2\sqrt{\alpha}}\left(\|\bm{W}\|^{2}_{F}+\alpha n\|\mkern 1.5mu\overline{\mkern-1.5mu\bm{H}\mkern-1.5mu}\mkern 1.5mu\|^{2}_{F}\right)=\|\mkern 1.5mu\overline{\mkern-1.5mu\bm{Z}\mkern-1.5mu}\mkern 1.5mu\|_{\ast}.$
This minimum is achieved only when $\bm{W}=\sqrt{\alpha n}\mkern 1.5mu\overline{\mkern-1.5mu\bm{H}\mkern-1.5mu}\mkern 1.5mu$, thus completing the proof. ∎
Appendix C Proof of Theorem 4.1
C.1 Proof of Theorem 4.1
In this section, we present the proof of Theorem 4.1 in Section 4, which we restate as follows.
Theorem C.1.
(Global Optimizer). Assume that the feature dimension $d$ is no less than the number of classes $K$, i.e., $d\geq K$, and the dataset is balanced. Then any global optimizer of $(\bm{W},\bm{H},\bm{b})$ of
$\min_{\bm{W},\bm{H},\bm{b}}\mathcal{L}(\bm{W},\bm{H},\bm{b}):=\frac{1}{N}{l_{CE}}(\bm{W}^{\top}\bm{H}+\bm{b}\mathbf{1}_{N}^{\top},\bm{Y}^{\delta})+\frac{\lambda_{W}}{2}\|\bm{W}\|^{2}_{F}+\frac{\lambda_{H}}{2}\|\bm{H}\|^{2}_{F}+\frac{\lambda_{b}}{2}\|\bm{b}\|^{2}$ (17)
satisfies the following properties:
The classification weight matrix $\bm{W}$ is given by
$\bm{W}=\left(\frac{\lambda_{H}n}{\lambda_{W}}\right)^{1/4}\sqrt{a^{\delta}}\bm{P}\left(K\bm{I}_{K}-\bm{1}_{K}\bm{1}_{K}^{\top}\right).$ (18)
The matrix of last-layer feature $\bm{H}$ can be represented as
$\bm{H}=\mkern 1.5mu\overline{\mkern-1.5mu\bm{H}\mkern-1.5mu}\mkern 1.5mu\bm{Y},\quad\mkern 1.5mu\overline{\mkern-1.5mu\bm{H}\mkern-1.5mu}\mkern 1.5mu=\left(\frac{\lambda_{W}}{\lambda_{H}n}\right)^{1/4}\sqrt{a^{\delta}}\bm{P}\left(K\bm{I}_{K}-\bm{1}_{K}\bm{1}_{K}^{\top}\right).$ (19)
The bias $\bm{b}$ is a zero vector, i.e. $\bm{b}=\bm{0}$.
Here, $\bm{P}\in\mathbb{R}^{d\times K}(d\geq K)$ is a partial orthogonal matrix such that $\bm{P}^{\top}\bm{P}=\bm{I}_{K}$ and $a^{\delta}$ satisfies:
  • (i)
    if $\sqrt{KN}\lambda_{Z}+\delta\geq 1$, then $a^{\delta}=0$;
  • (ii)
    if $\sqrt{KN}\lambda_{Z}+\delta<1$, then
    $a^{\delta}=\frac{1}{K}\log\left(\frac{K}{\sqrt{KN}\lambda_{Z}+\delta}-K+1)\right).$
with $\lambda_{Z}=\sqrt{\lambda_{W}\lambda_{H}}$.
Proof.
The main idea of proving Theorem 4.1 is first to connect the problem (17) to its corresponding convex counterpart. This allows us to derive the precise form of the global minimizer for the convex optimization problem. Subsequently, we can further characterize the specific structures of $\bm{W}$ and $\bm{H}$ based on the acquired global minimizer.
Connection of (17) to a Convex Problem. Let $\bm{Z}=\bm{W}^{\top}\bm{H}\in\mathbb{R}^{K\times N}$ represent the output logit matrix with $N=nK$ and $\alpha=\frac{\lambda_{H}}{\lambda_{W}}$. Utilizing Lemma B.3, we get:
$\min_{\bm{W}^{\top}\bm{H}=\bm{Z}}\lambda_{W}\|\bm{W}\|^{2}_{F}+\lambda_{H}\|\bm{H}\|^{2}_{F}$ $=\sqrt{\lambda_{W}\lambda_{H}}\min_{\bm{W}^{\top}\bm{H}=\bm{Z}}\frac{1}{\sqrt{\alpha}}\left(\|\bm{W}\|^{2}_{F}+\alpha\|\bm{H}\|^{2}_{F}\right)$ (20)
$=2\sqrt{\lambda_{W}\lambda_{H}}\|\bm{Z}\|_{\ast},$
where $\|\bm{Z}\|_{\ast}$ represent nuclear norm of $\bm{Z}$. Additionally from Lemma B.3, the minimum is attained only when $\bm{H}=\mkern 1.5mu\overline{\mkern-1.5mu\bm{H}\mkern-1.5mu}\mkern 1.5mu\bm{Y}$ and $\bm{W}=\sqrt{\alpha n}\mkern 1.5mu\overline{\mkern-1.5mu\bm{H}\mkern-1.5mu}\mkern 1.5mu=\sqrt{\frac{\lambda_{H}}{\lambda_{W}n}}\mkern 1.5mu\overline{\mkern-1.5mu\bm{H}\mkern-1.5mu}\mkern 1.5mu$.
Let $\lambda_{Z}:=\sqrt{\lambda_{W}\lambda_{H}}$, then Equation (17) becomes:
$\min_{\bm{Z},\bm{b}}\mathcal{L}(\bm{Z},\bm{b}):=\frac{1}{N}{l_{CE}}(\bm{Z}+\bm{b}\mathbf{1}_{N}^{\top},\bm{Y}^{\delta})+\lambda_{Z}\|\bm{Z}\|_{\ast}+\frac{\lambda_{b}}{2}\|\bm{b}\|^{2},$ (21)
which is a convex optimization problem.
Characterizing the Optimal Solution of (17) based on the Convex Program (21): We first derive the exact form of the global minimizer for the convex optimization problem (21). Particularly, we first establish that the predicted logit vectors within each class collapse to their sample means, i.e., $\bm{Z}={\mkern 1.5mu\overline{\mkern-1.5mu\bm{Z}\mkern-1.5mu}\mkern 1.5mu}\bm{Y}$ (Lemma C.3). Subsequently, we derive the closed-form solution of $\mkern 1.5mu\overline{\mkern-1.5mu\bm{Z}\mkern-1.5mu}\mkern 1.5mu$ in Lemma C.5.
Furthermore, from Lemma B.3, we establish that the minimum in (20) is attained only if $\bm{H}=\mkern 1.5mu\overline{\mkern-1.5mu\bm{H}\mkern-1.5mu}\mkern 1.5mu\bm{Y}$ and $\bm{W}=\sqrt{\frac{\lambda_{H}}{\lambda_{W}n}}\mkern 1.5mu\overline{\mkern-1.5mu\bm{H}\mkern-1.5mu}\mkern 1.5mu$. Since $\bm{Z}=\bm{W}^{\top}\bm{H}$, combining the above result with Lemma C.5 yields the global minimizer $(\bm{W},\bm{H},\bm{b})$ of (17), satisfying:
$\bm{W}=\left(\frac{\lambda_{H}n}{\lambda_{W}}\right)^{1/4}\sqrt{a^{\delta}}\bm{P}\left(K\bm{I}_{K}-\bm{1}_{K}\bm{1}_{K}^{\top}\right),$ (22)
$\bm{H}=\mkern 1.5mu\overline{\mkern-1.5mu\bm{H}\mkern-1.5mu}\mkern 1.5mu\bm{Y},\quad\mkern 1.5mu\overline{\mkern-1.5mu\bm{H}\mkern-1.5mu}\mkern 1.5mu=\left(\frac{\lambda_{W}}{\lambda_{H}n}\right)^{1/4}\sqrt{a^{\delta}}\bm{P}\left(K\bm{I}_{K}-\bm{1}_{K}\bm{1}_{K}^{\top}\right),$ (23)
and
$\bm{b}=\bm{0},$
where $\bm{P}\in\mathbb{R}^{d\times K}(d\geq K)$ is a partial orthogonal matrix such that $\bm{P}^{\top}\bm{P}=\bm{I}_{K}$ and $a^{\delta}$ is defined per the specifications in Lemma C.5.
Hence, we conclude the proof.
C.2 Supporting Lemma
Lemma C.2.
(Optimality Condition) The first-order optimality condition of $\mathcal{L}(\bm{Z},\bm{b})$ in Equation (21) is
$N^{-1}\left(\bm{Y}^{\delta}-\bm{P}\right)\in\lambda_{Z}\partial\|\bm{Z}\|_{\ast},\quad N^{-1}\left(\bm{Y}^{\delta}-\bm{P}\right)\mathbf{1}_{N}=\lambda_{b}\bm{b},$ (24)
where $\bm{P}$ is the prediction matrix defined as
$\bm{P}=\{\bm{p}_{ki}\}_{1\leq k\leq K,1\leq i\leq n}\in\mathbb{R}^{K\times N},\quad\bm{p}_{ki}:=\frac{\exp(\bm{z}_{ki}+\bm{b})}{\langle\exp(\bm{z}_{ki}+\bm{b}),\mathbf{1}_{K}\rangle},$ (25)
$\bm{Y}^{\delta}$ is the matrix of the smoothed soft target defined as
$\bm{Y}^{\delta}=(1-\delta)\bm{Y}+\frac{\delta}{K}\mathbf{1}_{K}\mathbf{1}_{N}^{\top},\quad\bm{Y}=[\bm{e}_{1}\mathbf{1}^{\top}_{n},\cdots,\bm{e}_{K}\mathbf{1}^{\top}_{n}]\in\mathbb{R}^{K\times N},$ (26)
and $\partial\|\bm{Z}\|_{\ast}$ represents the subdifferential of the nuclear norm of $\bm{Z}$.
Proof.
Consider
$\mathcal{L}(\bm{Z},\bm{b})=\frac{1}{N}{l_{CE}}(\bm{Z}+\bm{b}\mathbf{1}_{N}^{\top},\bm{Y}^{\delta})+\lambda_{Z}\|\bm{Z}\|_{\ast}+\frac{\lambda_{b}}{2}\|\bm{b}\|^{2}$
in (21). Define
$\phi(\bm{Z},\bm{b})=\frac{1}{N}l_{CE}(\bm{Z}+\bm{b}\mathbf{1}_{N}^{\top},\bm{Y}^{\delta})=\frac{1}{N}\sum_{k=1}^{K}\sum_{i=1}^{n}l_{CE}(\bm{z}_{ki}+\bm{b},\bm{y}^{\delta}_{k}),$ (27)
where $\bm{y}^{\delta}_{k}=(1-\delta)\bm{e}_{k}+\frac{\delta}{K}\mathbf{1}_{K}$ is the smoothed target. The gradient of $\phi$ is
$\frac{\partial\phi}{\partial\bm{z}_{ki}}=\frac{1}{N}\left(\bm{p}_{ki}-\bm{y}^{\delta}_{k}\right),\quad\frac{\partial\phi}{\partial\bm{b}}=\frac{1}{N}\sum_{k,i}\left(\bm{p}_{ki}-\bm{y}^{\delta}_{k}\right),$
whose matrix form is:
$\frac{\partial\phi}{\partial\bm{Z}}=\frac{1}{N}\left(\bm{P}-\bm{Y}^{\delta}\right),\quad\frac{\partial\phi}{\partial\bm{b}}=\frac{1}{N}\left(\bm{P}-\bm{Y}^{\delta}\right)\mathbf{1}_{N}.$
Hence the gradient (subgradient) of $\mathcal{L}$ is
$\frac{\partial\mathcal{L}}{\partial\bm{Z}}=N^{-1}(\bm{P}-\bm{Y}^{\delta})+\lambda_{Z}\partial\|\bm{Z}\|_{\ast},\quad\frac{\partial\mathcal{L}}{\partial\bm{b}}=N^{-1}(\bm{P}-\bm{Y}^{\delta})\mathbf{1}_{N}+\lambda_{b}\bm{b},$
where $\partial\|\bm{Z}\|_{\ast}$ is the subdifferential of the nuclear norm at $\bm{Z}$. Thus, $(\bm{Z},\bm{b})$ is a global minimizer of $\mathcal{L}$ if its gradient (subgradient) is equal to zero, i.e.
$N^{-1}(\bm{Y}^{\delta}-\bm{P})\in\lambda_{Z}\partial\|\bm{Z}\|_{\ast},\quad N^{-1}(\bm{Y}^{\delta}-\bm{P})\mathbf{1}_{N}=\lambda_{b}\bm{b}.$
Lemma C.3.
Assume that the number of classes $K$ is less than the feature dimension $d$, i.e., $K\leq d$, and the dataset is balanced. Then the prediction vectors, formulated as $\bm{z}_{ki}=f_{\Theta}(\bm{x}_{ki}),(1\leq k\leq K,1\leq i\leq n)$ within each class collapse to their sample means $\bar{\bm{z}}_{k}$:
$\bm{z}_{ki}=\bar{\bm{z}}_{k},\quad 1\leq i\leq n,$ (28)
In other words, the prediction matrix $\bm{Z}$ can be written as the following factorized form:
$\bm{Z}=\mkern 1.5mu\overline{\mkern-1.5mu\bm{Z}\mkern-1.5mu}\mkern 1.5mu\bm{Y}\in\mathbb{R}^{K\times N}$ (29)
where
$\mkern 1.5mu\overline{\mkern-1.5mu\bm{Z}\mkern-1.5mu}\mkern 1.5mu=[\bar{\bm{z}}_{1},\dots,\bar{\bm{z}}_{K}],\quad\bm{Y}=[\bm{e}_{1}\mathbf{1}_{n}^{\top},\cdots,\bm{e}_{K}\mathbf{1}_{n}^{\top}].$
Proof.
The proof follows from the convexity of the loss function. Recall that the loss function in (21) was given by
$\mathcal{L}(\bm{Z},\bm{b}):=\phi(\bm{Z},\bm{b})+\lambda_{Z}\|\bm{Z}\|_{\ast}+\frac{\lambda_{b}}{2}\|\bm{b}\|^{2},$
where $\phi(\bm{Z},\bm{b})$ is the Cross-Entropy (CE) or Label Smoothing (LS) loss (depending on the value of the smoothing parameter $\delta$) as defined in (27). Recalling that $\{\bm{z}_{ik}\}_{i=1}^{n}$ belong to the same class and $\bar{\bm{z}}_{k}=n^{-1}\sum_{i=1}^{n}\bm{z}_{ki}$ is their respective sample mean. From Jensen’s inequality, we have
$\phi(\bm{Z},\bm{b})$ $=\frac{1}{Kn}\sum_{k=1}^{K}\sum_{i=1}^{n}l_{CE}(\bm{z}_{ki}+b,\bm{y}_{k}^{\delta})$
$\geq\frac{1}{K}\sum_{k=1}^{K}l_{CE}\left(\frac{1}{n}\sum_{i=1}^{n}(\bm{z}_{ki}+b),\bm{y}_{k}^{\delta}\right)$
$=\frac{1}{K}\sum_{k=1}^{K}l_{CE}(\bar{\bm{z}}_{k}+\bm{b},\bm{y}_{k}^{\delta})$
where the inequality becomes equality only when $\bm{z}_{ki}=\bar{\bm{z}}_{k}$.
In the rest of the proof, we employ a permutation argument. Let $\bm{Z}_{l}=[\bm{z}_{1l},\cdots,\bm{z}_{Kl}]$ for $1\leq l\leq n$ and $\tilde{\bm{Z}}=[\bm{Z}_{1},\cdots,\bm{Z}_{n}]$. Let $\Gamma_{i}$ represent a distinct permutation of $n$. Consider $\tilde{\bm{Z}}_{\Gamma_{i}}=\tilde{\bm{Z}}\Pi_{\Gamma_{i}}$, where $\Pi_{\Gamma_{i}}$ is a permutation matrix rearranging $\tilde{\bm{Z}}$ so that the elements $\bm{Z}_{l}$ are ordered according to $\Gamma_{i}$. Then
$\|\tilde{\bm{Z}}_{\Gamma_{i}}\|_{\ast}=\|\tilde{\bm{Z}}\|_{\ast}=\|{\bm{Z}}\|_{\ast}.$
Since $\|\cdot\|_{\ast}$ is a convex function, we deduce
$\|{\bm{Z}}\|_{\ast}=\frac{1}{n!}\left(\sum_{i}\left\|\tilde{\bm{Z}}_{\Gamma_{i}}\right\|_{\ast}\right)\geq\left\|\frac{1}{n!}\sum_{i}\tilde{\bm{Z}}_{\Gamma_{i}}\right\|_{\ast},$
where the inequality becomes equality only when $\bm{Z}_{l}=\mkern 1.5mu\overline{\mkern-1.5mu\bm{Z}\mkern-1.5mu}\mkern 1.5mu(1\leq l\leq n)$ or equivalently $\bm{z}_{ki}=\bar{\bm{z}}_{k}$. As a result, it holds that
$\mathcal{L}(\bm{Z},\bm{b})$ $=\phi(\bm{Z},\bm{b})+\lambda_{Z}\|\bm{Z}\|_{\ast}+\frac{\lambda_{b}}{2}\|\bm{b}\|^{2}$
$\geq\frac{1}{K}\sum_{k=1}^{K}l_{CE}(\bar{z}_{k}+\bm{b},\bm{y}_{k}^{\delta})+\lambda_{Z}\|\mkern 1.5mu\overline{\mkern-1.5mu\bm{Z}\mkern-1.5mu}\mkern 1.5mu\|_{\ast}+\frac{\lambda_{b}}{2}\|\bm{b}\|^{2}$
$=\frac{1}{N}l_{CE}\left(\mkern 1.5mu\overline{\mkern-1.5mu\bm{Z}\mkern-1.5mu}\mkern 1.5muY+\bm{b}\mathbf{1}_{N}^{\top},\bm{Y}^{\delta}\right)+\lambda_{Z}\|\mkern 1.5mu\overline{\mkern-1.5mu\bm{Z}\mkern-1.5mu}\mkern 1.5mu\|_{\ast}+\frac{\lambda_{b}}{2}\|\bm{b}\|^{2}.$
The global optimality of $\mathcal{L}(\bm{Z},\bm{b})$ implies that $\bm{Z}=\mkern 1.5mu\overline{\mkern-1.5mu\bm{Z}\mkern-1.5mu}\mkern 1.5muY$, i.e., $\bm{z}_{ki}=\bar{\bm{z}}_{k}$ for $1\leq k\leq n$.
Lemma C.4.
Assume $\bm{z}_{ki}=\bar{\bm{z}}_{k}$, for $1\leq k\leq K,1\leq i\leq n$ and $\langle\bar{\bm{z}}_{k},\mathbf{1}_{K}\rangle=0$, then it holds that
$N^{-1}\left((1-\delta)\bm{I}_{K}+\frac{\delta}{K}\bm{J}_{K}-\mkern 1.5mu\overline{\mkern-1.5mu\bm{P}\mkern-1.5mu}\mkern 1.5mu\right)=\lambda_{Z}\left(n^{-1/2}\left[\left(\mkern 1.5mu\overline{\mkern-1.5mu\bm{Z}\mkern-1.5mu}\mkern 1.5mu\mkern 1.5mu\overline{\mkern-1.5mu\bm{Z}\mkern-1.5mu}\mkern 1.5mu^{\top}\right)^{\dagger}\right]^{1/2}\mkern 1.5mu\overline{\mkern-1.5mu\bm{Z}\mkern-1.5mu}\mkern 1.5mu+\mkern 1.5mu\overline{\mkern-1.5mu\bm{R}\mkern-1.5mu}\mkern 1.5mu\right),$
$\frac{n}{N}\left((1-\delta)\bm{I}_{K}+\frac{\delta}{K}\bm{J}_{K}-\mkern 1.5mu\overline{\mkern-1.5mu\bm{P}\mkern-1.5mu}\mkern 1.5mu\right)\mathbf{1}_{K}=\lambda_{b}\bm{b},$
[…]
In particular, if $\mkern 1.5mu\overline{\mkern-1.5mu\bm{Z}\mkern-1.5mu}\mkern 1.5mu$ is of rank $K-1$, then
$N^{-1}\left((1-\delta)\bm{I}_{K}+\frac{\delta}{K}\bm{J}_{K}-\mkern 1.5mu\overline{\mkern-1.5mu\bm{P}\mkern-1.5mu}\mkern 1.5mu\right)=\frac{\lambda_{Z}}{\sqrt{n}}\left(\left[\left(\mkern 1.5mu\overline{\mkern-1.5mu\bm{Z}\mkern-1.5mu}\mkern 1.5mu\mkern 1.5mu\overline{\mkern-1.5mu\bm{Z}\mkern-1.5mu}\mkern 1.5mu^{\top}\right)^{\dagger}\right]^{1/2}\mkern 1.5mu\overline{\mkern-1.5mu\bm{Z}\mkern-1.5mu}\mkern 1.5mu\right).$
Proof.
Under the assumption, $\bm{z}_{ki}=\bar{\bm{z}}_{k}$, $1\leq i\leq n$, and $\langle\bar{\bm{z}}_{k},\mathbf{1}_{K}\rangle=0$, we have $\bm{Z}=\mkern 1.5mu\overline{\mkern-1.5mu\bm{Z}\mkern-1.5mu}\mkern 1.5mu\bm{Y}$ and $\bm{P}=\mkern 1.5mu\overline{\mkern-1.5mu\bm{P}\mkern-1.5mu}\mkern 1.5mu\bm{Y}$, where $\mkern 1.5mu\overline{\mkern-1.5mu\bm{P}\mkern-1.5mu}\mkern 1.5mu$ is defined as
$\mkern 1.5mu\overline{\mkern-1.5mu\bm{P}\mkern-1.5mu}\mkern 1.5mu=[\bar{\bm{p}}_{1},\cdots,\bar{\bm{p}}_{K}]$
with $\bar{\bm{p}}_{k}$ as the probability vector w.r.t. the $\bar{\bm{z}}_{k}+\bm{b}$. Then the optimality condition in (24) reduces to
$N^{-1}\left((1-\delta)\bm{I}_{K}+\frac{\delta}{K}\bm{J}_{K}-\mkern 1.5mu\overline{\mkern-1.5mu\bm{P}\mkern-1.5mu}\mkern 1.5mu\right)\bm{Y}\in\lambda_{Z}\partial\|\bm{Z}\|_{\ast}.$ (30)
Let $\bm{Z}=\bm{U}\Sigma\bm{V}^{\top}$ be the SVD of $\bm{Z}$, then we have:
$\partial\|\bm{Z}\|_{\ast}=\{\bm{U}\bm{V}^{\top}+\bm{R}:\|\bm{R}\|\leq 1,\bm{U}^{\top}\bm{R}=0,\bm{R}\bm{V}=0\},$
where
$\bm{U}\bm{V}^{\top}=\left[\left(\bm{Z}\bm{Z}^{\top}\right)^{\dagger}\right]^{1/2}\bm{Z}.$
Since $\bm{Z}=\mkern 1.5mu\overline{\mkern-1.5mu\bm{Z}\mkern-1.5mu}\mkern 1.5muY$ and $\bm{Y}\bm{Y}^{\top}=n\bm{I}_{K}$, we further get:
$\bm{U}\bm{V}^{\top}=n^{-1/2}\left[\left(\mkern 1.5mu\overline{\mkern-1.5mu\bm{Z}\mkern-1.5mu}\mkern 1.5mu\mkern 1.5mu\overline{\mkern-1.5mu\bm{Z}\mkern-1.5mu}\mkern 1.5mu^{\top}\right)^{\dagger}\right]^{1/2}\mkern 1.5mu\overline{\mkern-1.5mu\bm{Z}\mkern-1.5mu}\mkern 1.5mu\bm{Y}.$
Then (30) is equivalent to:
$N^{-1}\left((1-\delta)\bm{I}_{K}+\frac{\delta}{K}\bm{J}_{K}-\mkern 1.5mu\overline{\mkern-1.5mu\bm{P}\mkern-1.5mu}\mkern 1.5mu\right)\bm{Y}=\lambda_{Z}\left(n^{-1/2}\left[\left(\mkern 1.5mu\overline{\mkern-1.5mu\bm{Z}\mkern-1.5mu}\mkern 1.5mu\mkern 1.5mu\overline{\mkern-1.5mu\bm{Z}\mkern-1.5mu}\mkern 1.5mu^{\top}\right)^{\dagger}\right]^{1/2}\mkern 1.5mu\overline{\mkern-1.5mu\bm{Z}\mkern-1.5mu}\mkern 1.5mu\bm{Y}+\bm{R}\right),$
where $\bm{R}$ is in the form of
$\bm{R}=\mkern 1.5mu\overline{\mkern-1.5mu\bm{R}\mkern-1.5mu}\mkern 1.5mu\bm{Y},\quad\mkern 1.5mu\overline{\mkern-1.5mu\bm{R}\mkern-1.5mu}\mkern 1.5mu=[\bar{\bm{r}}_{1},\cdots,\bar{\bm{r}}_{K}]$
such that
$\mkern 1.5mu\overline{\mkern-1.5mu\bm{R}\mkern-1.5mu}\mkern 1.5mu\bm{Y}\left(\mkern 1.5mu\overline{\mkern-1.5mu\bm{Z}\mkern-1.5mu}\mkern 1.5mu\bm{Y}\right)^{\top}=n\mkern 1.5mu\overline{\mkern-1.5mu\bm{R}\mkern-1.5mu}\mkern 1.5mu\mkern 1.5mu\overline{\mkern-1.5mu\bm{Z}\mkern-1.5mu}\mkern 1.5mu^{\top}=0,\quad\mkern 1.5mu\overline{\mkern-1.5mu\bm{Z}\mkern-1.5mu}\mkern 1.5mu^{\top}\mkern 1.5mu\overline{\mkern-1.5mu\bm{R}\mkern-1.5mu}\mkern 1.5mu=0,\quad\|n^{1/2}\mkern 1.5mu\overline{\mkern-1.5mu\bm{R}\mkern-1.5mu}\mkern 1.5mu\|\leq 1.$
This further leads to
$N^{-1}\left((1-\delta)\bm{I}_{K}+\frac{\delta}{K}\bm{J}_{K}-\mkern 1.5mu\overline{\mkern-1.5mu\bm{P}\mkern-1.5mu}\mkern 1.5mu\right)=\lambda_{Z}\left(n^{-1/2}\left[\left(\mkern 1.5mu\overline{\mkern-1.5mu\bm{Z}\mkern-1.5mu}\mkern 1.5mu\mkern 1.5mu\overline{\mkern-1.5mu\bm{Z}\mkern-1.5mu}\mkern 1.5mu^{\top}\right)^{\dagger}\right]^{1/2}\mkern 1.5mu\overline{\mkern-1.5mu\bm{Z}\mkern-1.5mu}\mkern 1.5mu+\mkern 1.5mu\overline{\mkern-1.5mu\bm{R}\mkern-1.5mu}\mkern 1.5mu\right)$
[…]
For $\bm{b}$, it is easy to see that the optimality condition in (24) reduces to
$\lambda_{b}\bm{b}=N^{-1}\left(\bm{Y}^{\delta}-\bm{P}\right)\mathbf{1}_{N}=\frac{n}{N}\left((1-\delta)\bm{I}_{K}+\frac{\delta}{K}\bm{J}_{K}-\mkern 1.5mu\overline{\mkern-1.5mu\bm{P}\mkern-1.5mu}\mkern 1.5mu\right)\mathbf{1}_{K}.$
Now, if $\mkern 1.5mu\overline{\mkern-1.5mu\bm{Z}\mkern-1.5mu}\mkern 1.5mu$ is of rank $K-1$, since $\mathbf{1}_{K}^{\top}\mkern 1.5mu\overline{\mkern-1.5mu\bm{Z}\mkern-1.5mu}\mkern 1.5mu=0$, the columns of $\mkern 1.5mu\overline{\mkern-1.5mu\bm{R}\mkern-1.5mu}\mkern 1.5mu$ are parallel to $\mathbf{1}_{K}$. Moreover, $\mkern 1.5mu\overline{\mkern-1.5mu\bm{P}\mkern-1.5mu}\mkern 1.5mu$ is a positive left stochastic matrix with $\mkern 1.5mu\overline{\mkern-1.5mu\bm{P}\mkern-1.5mu}\mkern 1.5mu^{\top}\mathbf{1}_{K}=\mathbf{1}_{K}$, and therefore $\bm{I}_{K}-\mkern 1.5mu\overline{\mkern-1.5mu\bm{P}\mkern-1.5mu}\mkern 1.5mu$ is also of rank $K-1$. The left stochasticity of $\mkern 1.5mu\overline{\mkern-1.5mu\bm{P}\mkern-1.5mu}\mkern 1.5mu$ further deduces that $(\bm{I}_{K}-\mkern 1.5mu\overline{\mkern-1.5mu\bm{P}\mkern-1.5mu}\mkern 1.5mu)^{T}\mathbf{1}_{K}=0$. In other words, $\mathbf{1}_{K}$ is in the left null space of $\bm{I}_{K}-\mkern 1.5mu\overline{\mkern-1.5mu\bm{P}\mkern-1.5mu}\mkern 1.5mu$, which in turn implies $\mkern 1.5mu\overline{\mkern-1.5mu\bm{R}\mkern-1.5mu}\mkern 1.5mu=0$. This leads to
$N^{-1}\left((1-\delta)\bm{I}_{K}+\frac{\delta}{K}\bm{J}_{K}-\mkern 1.5mu\overline{\mkern-1.5mu\bm{P}\mkern-1.5mu}\mkern 1.5mu\right)=\frac{\lambda_{Z}}{\sqrt{n}}\left(\left[\left(\mkern 1.5mu\overline{\mkern-1.5mu\bm{Z}\mkern-1.5mu}\mkern 1.5mu\mkern 1.5mu\overline{\mkern-1.5mu\bm{Z}\mkern-1.5mu}\mkern 1.5mu^{\top}\right)^{\dagger}\right]^{1/2}\mkern 1.5mu\overline{\mkern-1.5mu\bm{Z}\mkern-1.5mu}\mkern 1.5mu\right).$
Lemma C.5.
Assume that the number of classes $K$ is less than or equal to the feature dimension $d$, i.e., $K\leq d$, and the dataset is balanced. Then the global minimizer $(\bm{Z},\bm{b})$ of (21) satisfies the following properties:
$\bm{Z}=\mkern 1.5mu\overline{\mkern-1.5mu\bm{Z}\mkern-1.5mu}\mkern 1.5mu\bm{Y},\quad\mkern 1.5mu\overline{\mkern-1.5mu\bm{Z}\mkern-1.5mu}\mkern 1.5mu=a^{\delta}(K\bm{I}_{K}-\bm{J}_{K}),\quad\bm{b}=0.$ (31)
In Particular, we have
  • (i)
    if $\sqrt{KN}\lambda_{Z}+\delta\geq 1$, then $a^{\delta}=0$;
  • (ii)
    if $\sqrt{KN}\lambda_{Z}+\delta<1$, then
    $a^{\delta}=\frac{1}{K}\log\left(\frac{K}{\sqrt{KN}\lambda_{Z}+\delta}-K+1)\right)$
Proof.
From Lemma C.3, we have $\bm{Z}=\mkern 1.5mu\overline{\mkern-1.5mu\bm{Z}\mkern-1.5mu}\mkern 1.5muY$. Additionaly, combining Lemma 8 of (Zhou et al., 2022b), we have the global optimizer of (21) for different values of $\delta$ satisfies:
$\bm{Z}=a\left(K\bm{I}_{K}-\mathbf{1}_{K}\mathbf{1}_{K}^{\top}\right)\bm{Y},\quad\bm{b}=0.$
Equivalently, we have $\mkern 1.5mu\overline{\mkern-1.5mu\bm{Z}\mkern-1.5mu}\mkern 1.5mu=a\left(K\bm{I}_{K}-\mathbf{1}_{K}\mathbf{1}_{K}^{\top}\right)$. Further, we define $\mkern 1.5mu\overline{\mkern-1.5mu\bm{P}\mkern-1.5mu}\mkern 1.5mu=[\bar{\bm{p}}_{1},\cdots,\bar{\bm{p}}_{K}]$ with $\bar{\bm{p}}_{k}$ as the probability vector w.r.t. the logit $\bm{z}_{k}+\bm{b}$. Then,
$\mkern 1.5mu\overline{\mkern-1.5mu\bm{P}\mkern-1.5mu}\mkern 1.5mu=\frac{e^{-a}\bm{J}_{K}+e^{-a}(e^{aK}-1)\bm{I}_{K}}{e^{-a}(K-1+e^{aK})}=\frac{\bm{J}_{K}+(e^{aK}-1)\bm{I}_{K}}{K-1+e^{aK}},$ (32)
and
$(1-\delta)\bm{I}_{K}+\frac{\delta}{K}\bm{J}_{K}-\mkern 1.5mu\overline{\mkern-1.5mu\bm{P}\mkern-1.5mu}\mkern 1.5mu$ $=(1-\delta)\bm{I}_{K}+\frac{\delta}{K}\bm{J}_{K}-\frac{\bm{J}_{K}+(e^{aK}-1)\bm{I}_{K}}{K-1+e^{aK}}$
$=\left(\frac{K}{K-1+e^{aK}}-\delta\right)\left(\bm{I}_{K}-\frac{1}{K}\bm{J}_{K}\right)$
Recall that the first-order optimality condition as expressed in Lemma C.2 indicates that:
[…]
Suppose $a\neq 0$, then $\mkern 1.5mu\overline{\mkern-1.5mu\bm{Z}\mkern-1.5mu}\mkern 1.5mu$ has rank $K-1$ and we have $\mkern 1.5mu\overline{\mkern-1.5mu\bm{R}\mkern-1.5mu}\mkern 1.5mu=0$. Thus the above optimality condition implies:
$\left(\frac{K}{K-1+e^{aK}}-\delta\right)\left(\bm{I}_{K}-\bm{J}_{K}/K\right)=\text{sign}(a)\sqrt{KN}\lambda_{Z}(\bm{I}_{K}-\bm{J}_{K}/K),$
which is equivalent to
$\frac{K}{K-1+e^{aK}}-\delta=\text{sign}(a)\sqrt{NK}\lambda_{Z}.$
For $a>0$, the above equation has the following solution:
$a=\frac{1}{K}\log\left(\frac{K}{\sqrt{NK}\lambda_{Z}+\delta}-K+1\right)\quad\text{if}\quad\sqrt{NK}\lambda_{Z}+\delta<1.$
If $\sqrt{NK}\lambda_{Z}+\delta\geq 1$, then we select $a=0,$ and noting that in that case $\mkern 1.5mu\overline{\mkern-1.5mu\bm{P}\mkern-1.5mu}\mkern 1.5mu=K^{-1}\bm{J}_{K}$, we have that the optimality condition in (33) implies that $\mkern 1.5mu\overline{\mkern-1.5mu\bm{R}\mkern-1.5mu}\mkern 1.5mu$ satisfies
$(1-\delta)\bm{I}_{K}+\frac{\delta-1}{K}\bm{J}_{K}=N\lambda_{Z}\mkern 1.5mu\overline{\mkern-1.5mu\bm{R}\mkern-1.5mu}\mkern 1.5mu.$
We get
$\mkern 1.5mu\overline{\mkern-1.5mu\bm{R}\mkern-1.5mu}\mkern 1.5mu=\frac{1}{N\lambda_{Z}}\left[(1-\delta)\bm{I}_{K}+\frac{\delta-1}{K}\bm{J}_{K}\right],$
where $\|\mkern 1.5mu\overline{\mkern-1.5mu\bm{R}\mkern-1.5mu}\mkern 1.5mu\|\leq n^{-1/2}$ meets the requirement of the optimality condition.
Appendix D Proof of Theorem 4.2
In this section of the appendices, we prove Theorem 4.2.
Proof.
To provide insight into the accelerated convergence of the model under label smoothing loss, we examine the Hessian matrix of the empirical loss function, defined as follows:
$\min_{\bm{W},\bm{H}}\phi(\bm{W},\bm{H}):=\frac{1}{N}{l_{CE}}(\bm{W}^{\top}\bm{H},\bm{Y}^{\delta}).$ (35)
Under commonly used deep learning frameworks, the model parameters are updated iteratively, and it is common and often practical to analyze the Hessian matrix with respect to $\bm{H}$, $\bm{W}$ individually rather than considering the full Hessian. Particularly, we demonstrate that the Hessian of the empirical loss concerning ${\bm{W}}$ when ${\bm{H}}$ is fixed and the Hessian w.r.t. ${\bm{H}}$ when ${\bm{W}}$ is fixed are semi-definite at the global optimizer under both cross-entropy and label smoothing losses. Furthermore, we establish that the condition numbers of these Hessian matrices are notably lower under the LS loss in comparison to the CE loss.
Hessian matrix with respect to $\bm{Z}$. Let $\bm{Z}=\bm{W}^{\top}\bm{H}\in\mathbb{R}^{K\times N}$ represent the prediction logit matrix. In the proof of Lemma C.2, we obtained the first-order partial derivatives of the loss $\phi$:
$\frac{\partial\phi}{\partial\bm{z}_{ki}}=\frac{1}{N}\left(\bm{p}_{ki}-\bm{y}^{\delta}_{k}\right),\quad\frac{\partial\phi}{\partial\bm{b}}=\frac{1}{N}\sum_{k,i}\left(\bm{p}_{ki}-\bm{y}^{\delta}_{k}\right),$
where $\bm{p}_{ki}$ is the prediction for the $i$-th sample that belongs to the $k$-th class and $\bm{y}^{\delta}_{k}$ is the soft label for class $k$ with a smoothing parameter $\delta$. These partial derivatives lead to the corresponding second-order partial derivatives of the loss $\phi$:
$\frac{\partial^{2}\phi}{\partial\bm{z}^{2}_{ki}}=\frac{1}{N}\left(\text{diag}(\bm{p}_{ki})-\bm{p}_{ki}\bm{p}_{ki}^{\top}\right),\quad\frac{\partial^{2}\phi}{\partial\bm{z}_{ki}\partial\bm{z}_{k^{\prime}i^{\prime}}}=0,\forall(k,i)\neq(k^{\prime},i^{\prime}),$
$\ \frac{\partial^{2}\phi}{\partial\bm{b}^{2}}=\frac{1}{N}\sum_{k,i}\left(\text{diag}(\bm{p}_{ki})-\bm{p}_{ki}\bm{p}_{ki}^{\top}\right),\quad\frac{\partial^{2}\phi}{\partial\bm{z}_{ki}\partial\bm{b}}=\frac{1}{N}\left(\text{diag}(\bm{p}_{ki})-\bm{p}_{ki}\bm{p}_{ki}^{\top}\right).$
From Theorem 4.1, we have the global optimizer satisfies $\bm{z}^{2}_{ki}=\bar{\bm{z}}_{k}$ and $\bm{p}^{2}_{ki}=\bar{\bm{p}}_{k},1\leq k\leq K,1\leq i\leq n$. Consequently, it follows that $\frac{\partial^{2}\phi}{\partial\bm{z}^{2}_{ki}}=\frac{\partial^{2}\phi}{\partial\bm{z}^{2}_{ki^{\prime}}}(\forall 1\leq i,i^{\prime}\leq n)$. To simply the notation, we denote
$\bm{D}_{k}=\text{diag}(\bar{\bm{p}}_{k})-\bar{\bm{p}}_{k}\bar{\bm{p}}_{k}^{\top}.$ (36)
Then we have $\frac{\partial^{2}\phi}{\partial\bm{z}^{2}_{ki}}=\bm{D}_{k}/N$.
The closed-form solution for $\mkern 1.5mu\overline{\mkern-1.5mu\bm{P}\mkern-1.5mu}\mkern 1.5mu$ in (32) yields the expression:
$\bar{\bm{p}}_{k}=\frac{1}{K-1+e^{a^{\delta}K}}\left((e^{a^{\delta}K}-1)\bm{e}_{k}+\mathbf{1}_{K}\right).$ (37)
To simplify the notation, we denote
$p_{t}=e^{a^{\delta}K}/(K-1+e^{a^{\delta}K}),\quad p_{n}=1/(K-1+e^{a^{\delta}K}),$ (38)
where $p_{t}$ ($p_{n}$) are the predicted probability for the target class (non-target) class at the global optimal solution. Under this notation, we have $p_{t}+(K-1)p_{n}=1$ and $\bar{\bm{p}}_{k}=(p_{t}-p_{n})\bm{e}_{k}+p_{n}\mathbf{1}_{K}$.
For any $1\leq k\leq K$, $\bm{p}_{k}\bm{p}_{k}^{\top}$ is a positive matrix and the associated Laplacian $\bm{D}_{k}$ is positive-semidefinite with all eigenvalues non negative. The smallest eigenvalue of the Laplacian matrix $\bm{D}_{k}$ is $\sigma_{1}=0$ with the corresponding eigenvector $\bm{v}_{1}=\mathbf{1}_{K}/\|\mathbf{1}_{K}\|$.
Define
$\bm{v}_{2}=(K\bm{e}_{1}-\mathbf{1}_{K})/\|K\bm{e}_{1}-\mathbf{1}_{K}\|,$
then we have
$\bm{D}_{k}\bm{v}_{2}$ $=\text{diag}\left(\bar{\bm{p}}_{k}\right)\bm{v}_{2}-\bar{\bm{p}}_{k}\bar{\bm{p}}_{k}^{\top}\bm{v}_{2}$
$=Kp_{t}p_{n}\bm{v}_{2},$
from which we get $Kp_{t}p_{n}$ is an eigenvalue of $\bm{D}_{k}$ with the corresponding eigenvector $\bm{v}_{2}$.
Consider the outer-product matrix $\bm{p}_{k}\bm{p}_{k}^{\top}$, its largest eigenvalue is $\|\bm{p}_{k}\|^{2}$ with the eigenvector $\bm{p}_{k}/\|\bm{p}_{k}\|$. Its null space is of dimension $K-1$. Particularly, we can find a set of the basis vectors for $\text{Null}(\bm{p}_{k}\bm{p}_{k}^{\top})$ as follows:
$\{\bm{o}_{1},\cdots,\bm{o}_{K-1}:\text{span}(\bm{o}_{1},\bm{p}_{k})=\text{span}(\bm{e}_{k},\bm{p}_{k})\},$ (39)
which means the vectors $\bm{o}_{1}$ and $\bm{p}_{k}$ span the same 2D space as $\bm{e}_{k}$ and $\bm{p}_{k}$.
Note that
$\text{span}(\bm{e}_{k},\bm{p}_{k})=\text{span}(\bm{v}_{1},\bm{v}_{2}),$
where $\bm{v}_{1},\bm{v}_{2}$ are the eigenvector for $\bm{D}_{k}$.
Then we can show that $\bm{o}_{2}\cdots,\bm{o}_{K-1}$ are eigenvectors of $\bm{D}_{k}$ with the corresponding eigenvalue as $\sigma_{3}=p_{n}$. Particularly, with $\bm{o}_{l}$ orthonormal to both $\bar{\bm{p}}_{k}$ and $\bm{e}_{k}$, we have
$\bm{D}_{k}\bm{o}_{l}$ $=\text{diag}\left(\bar{\bm{p}}_{k}\right)\bm{o}_{l}-\bar{\bm{p}}_{k}\bar{\bm{p}}_{k}^{\top}\bm{o}_{l}$
$=\text{diag}\left(\bar{\bm{p}}_{k}\right)\bm{o}_{l}$
$=p_{n}\bm{o}_{l},$
for $\bm{o}_{l}(2\leq l\leq K-1)$ defined in (39).
To summarize the unique eigenvalues of the matrix $\bm{D}$ are
$\sigma_{1}=0,\quad\sigma_{2}=p_{n},\quad\sigma_{3}=Kp_{t}p_{n}$ (40)
with $p_{t}$ and $p_{n}$ defined in (38).
Hessian matrix with respect to $\bm{H}$. The gradient of $\phi$ with respect to $\bm{h}_{ki}$ is
$\frac{\partial\phi}{\partial\bm{h}_{ki}}=\frac{\partial\bm{z}_{ki}}{\partial\bm{h}_{ki}}\frac{\partial\phi}{\partial\bm{z}_{ki}}=\bm{W}\frac{\partial\phi}{\partial\bm{z}_{ki}}.$
Further, we can easily get the corresponding Hessian:
$\frac{\partial^{2}\phi}{\partial\bm{h}^{2}_{ki}}=\bm{W}\frac{\partial^{2}\phi}{\partial\bm{z}^{2}_{ki}}\bm{W}^{\top}=\frac{1}{N}\bm{W}\bm{D}_{k}\bm{W}^{\top},\quad\frac{\partial^{2}\phi}{\partial\bm{h}_{ki}\partial\bm{h}_{k^{\prime}i^{\prime}}}=0,$ (41)
where $\bm{D}_{k}$ is the Laplacian matrix as defined in (36).
From Theorem 4.1, we have
$\bm{W}=\left(\frac{\lambda_{H}n}{\lambda_{W}}\right)^{1/4}\sqrt{a^{\delta}}\bm{P}\left(K\bm{I}_{K}-\bm{J}_{K}\right).$
For simplicity, we denote $a^{\delta}_{W}=K(\lambda_{H}n/\lambda_{W})^{1/4}\sqrt{a^{\delta}}$, then $\bm{W}=a^{\delta}_{W}\bm{P}(\bm{I}_{K}-\bm{J}_{K}/K)$. Under this notation, we have
$\frac{1}{N}\bm{W}\bm{D}_{k}\bm{W}^{\top}$ $=\frac{(a^{\delta}_{W})^{2}}{N}\bm{P}(\bm{I}_{K}-\bm{J}_{K}/K)\bm{D}_{k}(\bm{I}_{K}-\bm{J}_{K}/K)\bm{P}^{\top}$ (42)
$=\frac{(a^{\delta}_{W})^{2}}{N}\bm{P}\bm{D}_{k}\bm{P}^{\top}.$ (43)
Note that $\bm{P}\in\mathbb{R}^{d\times K},(d>K)$ is a partial orthogonal matrix. Given the eigenvalues of $\bm{D}_{k}$ in (40), the eigenvalues for (42) are:
$\lambda_{1}=0,\quad\lambda_{2}=\frac{(a^{\delta}_{W})^{2}}{N}p_{n},\quad\lambda_{3}=\frac{(a^{\delta}_{W})^{2}}{N}Kp_{t}p_{n}$ (44)
with the corresponding multiplicities $m(\lambda_{1})=1+d-K$, $m(\lambda_{2})=K-2$, $m(\lambda_{3})=1$.
From (47), the hessian of $\phi$ with respect to $\bm{h}=\text{vec}(\bm{H})$ is a block diagonal matrix which can be expressed as
$\frac{\partial^{2}\phi}{\partial\bm{h}^{2}}=\frac{1}{N}\text{blkdiag}(\bm{W}\bm{D}_{1}\bm{W}^{\top},\cdots,\bm{W}\bm{D}_{1}\bm{W}^{\top},\cdots,\bm{W}\bm{D}_{K}\bm{W}^{\top},\cdots,\bm{W}\bm{D}_{K}\bm{W}^{\top}).$ (45)
The unique eigenvalues of the Hessian matrix $\frac{\partial^{2}\phi}{\partial\bm{h}^{2}}$ are the same as provided in (44). Given that the Hessian matrix contains a zero eigenvalue, our analysis centers on its condition number within the non-zero eigenvalue space. This is calculated as follows:
$\kappa(\nabla^{2}_{\bm{h}}\phi)=\lambda_{3}/\lambda_{2}=Kp_{t}$ (46)
Considering the formula for $p_{t}$ given in (38), an increase in $\delta$ leads to a decrease in $p_{t}$, subsequently resulting in a smaller condition number.
Hessian matrix with respect to $\bm{W}$. The gradient of $\phi$ with respect to $\bm{w}_{l}(l=1,\cdots,K)$ is
$\frac{\partial\phi}{\partial\bm{w}_{l}}=\frac{\partial\bm{z}_{ki}}{\partial\bm{w}_{l}}\frac{\partial\phi}{\partial\bm{z}_{ki}}.$
Further, we get the corresponding Hessian:
$\frac{\partial^{2}\phi}{\partial\bm{w}_{l}\partial\bm{w}_{l^{\prime}}}$ $=\sum_{k,i}\frac{\partial\bm{z}_{ki}}{\partial\bm{w}_{l^{\prime}}}\frac{\partial^{2}\phi}{\partial\bm{z}_{ki}^{2}}\left(\frac{\partial\bm{z}_{ki}}{\partial\bm{w}_{l}}\right)^{\top}$
$=\sum_{k,i}\bm{h}_{ki}\bm{e}_{l^{\prime}}^{\top}\frac{\partial^{2}\phi}{\partial\bm{z}_{ki}^{2}}\bm{e}_{l}\bm{h}_{ki}^{\top}$
$=\sum_{k,i}\bm{D}_{k}(l^{\prime},l)\bm{h}_{ki}\bm{h}_{ki}^{\top}$
$=n\sum_{k}\bm{D}_{k}(l^{\prime},l)\bar{\bm{h}}_{k}\bar{\bm{h}}_{k}^{\top}$
where $\bm{D}_{k}(l^{\prime},l)$ is $(l^{\prime},l)$-th element in $\bm{D}_{k}$ and $\bar{\bm{h}}_{k}$ is the $k$-th column vector in $\mkern 1.5mu\overline{\mkern-1.5mu\bm{H}\mkern-1.5mu}\mkern 1.5mu$ as defined in 19. From the defination of $\bm{D}_{k}$ in (36), we have
[…]
where $\bar{\bm{p}}_{l}$ is the average prediction vector as defined in (37). Particularly, the $l$-th element of $\bar{\bm{p}}_{l}$ equals $p_{t}$ and the others equal $p_{n}$.
From (19), we have
$\mkern 1.5mu\overline{\mkern-1.5mu\bm{H}\mkern-1.5mu}\mkern 1.5mu=\left(\frac{\lambda_{W}}{\lambda_{H}n}\right)^{1/4}\sqrt{a^{\delta}}\bm{P}\left(K\bm{I}_{K}-{\bm{J}_{K}}\right).$
Hence, the Hessian of $\phi$ w.r.t. the $\bm{w}=\text{vec}(\bm{W})$ can be written as:
$\frac{\partial^{2}\phi}{\partial\bm{w}^{2}}=nK^{2}\left(\frac{\lambda_{W}}{\lambda_{H}n}\right)^{1/2}a^{\delta}\bm{D}_{P}\bm{D}_{\Pi}\bm{S}\bm{D}_{\Pi}\bm{D}_{P}^{\top}$ (47)
where
$\bm{D}_{P}=\text{blkdiag}({\bm{P}},\cdots,{\bm{P}})\in\mathbb{R}^{Kd\times K^{2}},$ (48)
$\bm{D}_{\Pi}=\text{blkdiag}(\bm{\Pi},\cdots,\bm{\Pi})\in\mathbb{R}^{K^{2}\times K^{2}},\quad\bm{\Pi}=\bm{I}_{K}-\frac{\bm{1}_{K}\bm{1}_{K}^{\top}}{K}$ (49)
and
$\bm{S}=\begin{bmatrix}\Lambda_{1}&&\\ &\ddots&\\ &&\Lambda_{K}\end{bmatrix}-\begin{bmatrix}\Lambda_{1}\\ \vdots\\ \Lambda_{K}\end{bmatrix}\begin{bmatrix}\Lambda_{1}&\cdots&\Lambda_{K}\end{bmatrix}$ (50)
with $\Lambda_{l}=\text{diag}\left(\bar{\bm{p}}_{l}\right)$.
Since $\bm{P}$ is a partial orthogonal matrix, the condition number of (47) is the same as the condition number of the following matrix:
$\bm{B}=\bm{D}_{\Pi}\bm{S}\bm{D}_{\Pi}.$ (51)
Subsequently, we proceed to derive the eigenvalues for the above matrix $\bm{B}$.
First, considering an eigenvector $\bm{v}$ of $\bm{B}$, let $\Delta$ be a $K\times K$ matrix such that $\bm{v}=\text{vec}(\Delta)$. Considering any $\Delta$ satisfying $\Delta\bm{\Pi}=0$ or $\Delta^{\top}\bm{\Pi}=0$, for the corresponding $\bm{v}=\text{vec}(\Delta)$ we get $\bm{D}_{\Pi}\bm{v}=0$ and $\bm{v}^{\top}\bm{D}_{\Pi}=0$, wich further yields $\bm{B}\bm{v}=0$. Since $\text{rank}(\bm{\Pi})=K-1$, it follows that $\lambda_{1}=0$ is an eigenvalue of $\bm{B}$ with multiplicity $12K-$ 1.
On the other hand, consider any $\Delta$ that satisfies the conditions:
$\text{diag}(\Delta)=0,\quad\Delta\bm{1}_{K}=\Delta^{\top}\bm{1}_{K}=0.$ (52)
It is noteworthy that from $\Delta\bm{1}_{K}=\Delta^{\top}\bm{1}_{K}=0$, the corresponding vector $\bm{v}$ satisfies $\bm{D}_{\Pi}\bm{v}=\bm{v}$. If $\Delta$ further satisfies $\text{diag}(\Delta)=0$ then $\bm{S}\bm{v}=p_{n}\bm{v}$. Consequently, we deduce that
$\bm{B}\bm{v}=\bm{D}_{\Pi}\bm{S}\bm{D}_{\Pi}\bm{v}=\bm{D}_{\Pi}\bm{S}\bm{v}=\bm{D}_{\Pi}p_{n}\bm{v}=p_{n}\bm{v}$
We can conclude that $p_{n}$ is an eigenvalue of $\bm{B}$ with multiplicity $K^{2}-K-(2K-1)=K^{2}-3K+1$.
Next, let us consider the matrix $\Delta=\bm{\Pi}\text{diag}(\bm{u})\bm{\Pi}$, where $\bm{u}$ is a vector satisfying $\bm{u}^{\top}\bm{1}_{K}=0$. The vectorized form of $\Delta$ can be denoted as
$\bm{v}=\text{vec}(\Delta)=\begin{bmatrix}\bm{\Pi}^{\top}\text{diag}(\bm{u})\bm{\Pi}_{1}\\ \bm{\Pi}^{\top}\text{diag}(\bm{u})\bm{\Pi}_{2}\\ \vdots\\ \bm{\Pi}^{\top}\text{diag}(\bm{u})\bm{\Pi}_{K}\\ \end{bmatrix},$
where $\bm{\Pi}_{k}$ represents the $k$-th column of the matrix $\bm{\Pi}$. Due to the properties $\bm{\Pi}=\bm{\Pi}^{\top}$ and $\bm{\Pi}^{2}=\bm{\Pi}$, it follows that $\bm{D}_{\Pi}\bm{v}=\bm{v}$. Additionally, given $\bm{u}^{\top}\bm{1}_{K}=0$, we obtain $\bm{S}\bm{v}=\lambda_{2}\bm{v}$ with $\lambda_{2}=(1-p_{t}+p_{n})\cdot\frac{p_{n}+(K-1)p_{t}}{K}$. Consequently, we conclude that $\lambda_{2}=(1-p_{t}+p_{n})\cdot\frac{p_{n}+(K-1)p_{t}}{K}$ is an eigenvalue of $\bm{B}$ with a multiplicity equal to the degrees of freedom of $\bm{u}$, which is $K-1$.
Lastly, considering $\bm{v}=\text{vec}(\bm{\Pi})$, we observe that $\bm{D}_{\Pi}\bm{v}=\bm{v}$. Further, we have $\bm{S}\bm{v}=Kp_{t}p_{n}\bm{v}$. Hence, we obtain
$\bm{B}\bm{v}=\bm{D}_{\Pi}\bm{S}\bm{D}_{\Pi}\bm{v}=\bm{D}_{\Pi}\bm{S}\bm{v}=Kp_{t}p_{n}\bm{v}.$
Therefore, $\bm{v}=\text{vec}(\bm{\Pi})$ is an eigenvector of $\bm{B}$ with eigenvalue $\lambda_{3}=Kp_{t}p_{n}$.
In summary, we identify the distinct eigenvalues of the matrix $\bm{B}$ as defined in (51) as follows:
$\lambda_{0}=0,\quad\lambda_{1}=p_{n},\quad\lambda_{2}=(1-p_{t}+p_{n})\cdot\frac{p_{n}+(K-1)p_{t}}{K},\quad\lambda_{3}=Kp_{n}p_{t}$ (53)
with the corresponding multiplicities $12K-$ 1, $K^{2}-3K+1$, $K-1$, and $1$, respectively.
Since the matrix $\bm{B}$ has zero eigenvalues, we consider its condition number within the non-zero eigenvalue subspace, which is given by $\lambda_{3}/\lambda_{1}=Kp_{t}$. Consequently, we straightforwardly determine the condition number of the Hessian $\frac{\partial^{2}\phi}{\partial\bm{w}^{2}}$ as defined in (47) is
$\kappa(\nabla^{2}_{\bm{w}}\phi)=Kp_{t}.$ (54)
From the formula of $p_{t}$ given in (38), it is evident that increasing $\delta$ leads to a decrease of $p_{t}$, and consequently a reduction in the condition number.
By combining (46) and (54), we demonstrate that the Hessian matrices, $\nabla^{2}_{\bm{W}}\phi(\bm{W},\bm{H})$ and $\nabla^{2}_{\bm{H}}\phi(\bm{W},\bm{H})$, are positive semi-definite at the global optimizer. Moreover, the condition numbers of these Hessian matrices are notably lower under the label smoothing loss (with $10<\delta<$ 1) compared to the cross-entropy loss (with $\delta=0$) . This observation suggests that the optimization landscape of the empirical loss function is better conditioned around its global minimizer under label smoothing loss, thus completing our proof.
In addition, we emphasize that the above conclusion holds for an appropriate range of $\delta$ which is not close to $1$. As $\delta$ approaches $1$, the model does not provide a meaningful supervising signal. Consequently, considering the convergence rate under such conditions becomes impractical and does not yield meaningful insights.